### v2

# Dual Quaternion Laplacian Matrix and Formation Control

Abstract
The dual quaternion Laplacian matrix of desired relative configurations in multi-agent formation control is similar to the classical unweighted Laplacian matrix via a dual quaternion diagonal unitary matrix. Its eigenvalues are all positive numbers except one zero eigenvalue. A unit dual quaternion vector is a desired formation vector if and only if it is in the null space of this dual quaternion Laplacian matrix. We study a control law based upon dual quaternion Laplacian. We extend our discussion to directed graphs. We also show that pairwise asymptotical stability can be reduced to rank-one asymptotical stability. Key words. Dual quaternion Laplacian, formation control, relative configuration, stability, eigenvalues.
1 Introduction
The multi-agent formation control problem considers the formation control problem of $n$ rigid bodies in the 3-D space. It is a very important research problem in robotics research. Examples of these $n$ rigid bodies can be autonomous mobile robots, or unmanned aerial vehicles (UAVs), or autonomous underwater vehicles (AUVs), or small satellites. As the unit dual quaternion is a unified mathematical tool to describe the rigid body movement in the 3-D space, dual quaternion is used as a mathematical tool to study the formation control problem [14, 15]. Wang, Yu and Zheng [16] studied the application of dual quaternion matrices in the multi-agent formation control problem. In 2023, Qi, Wang and Luo [13] showed that the relative configuration adjacency matrix and the logarithm adjacency matrix in the multi-agent formation control problem are dual quaternion Hermitian matrices, which have good spectral properties as shown in [12]. They also introduced dual quaternion Laplacian matrices and studied their role in formation control. In this paper, we further study the applications of dual quaternion matrices in multi-agent formation control. In the next section, we review some basic definitions and properties of dual quaternions and dual quaternion matrices. A main aim of formation control is to make the relative configuration formation of these $n$ rigid bodies asymptotically converge to some desired relative configuration formation. Assume that any two of these $n$ rigid bodies can sense each other. It was called in [15] as pairwise asymptotical stability if all the $n^{2}$ pairwise relative configurations asymptotically converge to the desired relative configurations, as used as the goal of the formation control. In Section 3, we show that this is not necessary. It only needs some $n-1$ relative configurations asymptotically converge to the corresponding desired relative configurations. This simplifies the task of the formation control. In the multi-agent formation control, the Laplacian matrix of the mutual visibility graph plays a key role. In [8], Lin, Wang, Han and Fu presented a systematic approach based on complex Laplacian for formation control. They introduced the notion of similar formation and showed that all similar formations are those lie in the null space of a complex Laplacian satisfying certain rank conditions. This approach was further studied in [9, 7, 3, 4, 18]. The null space of the complex Laplacian given in [8] is somewhat complicated. The complex Laplacian given there is not Hermitian. This also increases the complexity of that approach. On the other hand, the dual quaternion is a unified mathematical tool to describe the rigid body movement in the 3-D space [15]. In Section 4 of this paper, we show that the dual quaternion Laplacian introduced in [13] is simpler and more natural. We show that the dual quaternion Laplacian of the desired relative configurations is similar to the real unweighted Laplacian, the eigenvalues of the dual quaternion Laplacian are all positive real numbers and one zero, and a unit dual quaternion vector is a desired formation vector if and only if it is an eigenvector of the zero eigenvalue of the dual quaternion Laplacian. In Section 5, we study a control law based upon the dual quaternion Laplacian for bidirectional and connected graphs. Then in Section 6, we extend our discussion to directed graphs, where the dual quaternion Laplacian is no longer Hermitian.
2 Dual Quaternions and Dual Quaternion Matrices
2.1 Quaternions and Dual Quaternions
Denote the field of real numbers by $\mathds{R}$, and the space of three-dimensional real vectors by $\mathbb{V}$. Use small letters with overline arrows, such as $\vec{p},\vec{q}$, to denote three-dimensional vectors. For $\vec{p}\in\mathbb{V}$, we denote that $\vec{p}=[p_{1},p_{2},p_{3}]$, but we still regard $\vec{p}$ as a column vector and its transpose $\vec{p}^{\top}$ a row vector. In particular, denote $\vec{0}:=[0,0,0]$. A quaternion $\tilde{q}=[q_{0},\vec{q}\,]=[q_{0},q_{1},q_{2},q_{3}]$ is a real four-dimensional vector, where $q_{0}\in\mathds{R}$ and $\vec{q}=[q_{1},q_{2},q_{3}]\in\mathbb{V}$. We use a tilde symbol to distinguish a quaternion. See [14]. Denote the set of all quaternions by $\mathbb{Q}$. Let $\tilde{0}:=[0,0,0,0]\in\mathbb{Q}$ be the zero element of $\mathbb{Q}$. Let $\tilde{p}=[p_{0},\vec{p}\,],\tilde{q}=[q_{0},\vec{q}\,]\in\mathbb{Q}$, where $\vec{p}=[p_{1},p_{2},p_{3}],\vec{q}=[q_{1},q_{2},q_{3}]\in\mathbb{V}$. The sum of $\tilde{p}$ and $\tilde{q}$ is defined as
$\tilde{p}+\tilde{q}=[p_{0}+q_{0},\vec{p}+\vec{q}\,].$
The product of $\tilde{p}$ and $\tilde{q}$ is defined by
$\tilde{p}\tilde{q}=[p_{0}q_{0}-\vec{p}\cdot\vec{q},p_{0}\vec{q}+q_{0}\vec{p}+\vec{p}\times\vec{q}\,],$
where $\vec{p}\cdot\vec{q}$ is the dot product, i.e., inner product of $\vec{p}$ and $\vec{q}$, with
$\vec{p}\cdot\vec{q}\equiv\vec{p}^{\top}\vec{q}=p_{1}q_{1}+p_{2}q_{2}+p_{3}q_{3},$
and $\vec{p}\times\vec{q}$ is the cross product of $\vec{p}$ and $\vec{q}$, with
$\vec{p}\times\vec{q}=[p_{2}q_{3}-p_{3}q_{2},-p_{1}q_{3}+p_{3}q_{1},p_{1}q_{2}-p_{2}q_{1}]=-\vec{q}\times\vec{p}.$
Thus, in general, $\tilde{p}\tilde{q}\not=\tilde{q}\tilde{p}$, and we have $\tilde{p}\tilde{q}=\tilde{q}\tilde{p}$ if and only if $\vec{p}\times\vec{q}=\vec{0}$, i.e., either $\vec{p}=\vec{0}$ or $\vec{q}=\vec{0}$, or $\vec{p}=\alpha\vec{q}$ for some real number $\alpha$. Though the quaternion multiplication is not commutative, but it is associative, i.e., for any $\tilde{p},\tilde{q},\tilde{s}\in\mathbb{Q}$, we have
$(\tilde{p}\tilde{q})\tilde{s}=\tilde{p}(\tilde{q}\tilde{s}).$
The conjugate of a quaternion $\tilde{q}=[q_{0},q_{1},q_{2},q_{3}]\in\mathbb{Q}$ is defined as $\tilde{q}^{*}=[q_{0},-q_{1},-q_{2},-q_{3}]$. Let $\tilde{1}:=[1,0,0,0]\in\mathbb{Q}$. Then for any $\tilde{q}\in\mathbb{Q}$, we have $\tilde{q}\tilde{1}=\tilde{1}\tilde{q}=\tilde{q}$, i.e., $\tilde{1}$ is the idenity element of $\mathbb{Q}$. For any $\tilde{p},\tilde{q}\in\mathbb{Q}$, we have
$(\tilde{p}\tilde{q})^{*}=\tilde{q}^{*}\tilde{p}^{*}.$
Suppose that $\tilde{p},\tilde{q}\in\mathbb{Q}$, $\tilde{p}\tilde{q}=\tilde{q}\tilde{p}=\tilde{1}$. Then we say that $\tilde{p}$ is invertible and its inverse is $\tilde{p}^{-1}=\tilde{q}$. Suppose that $\tilde{q}=[q_{0},q_{1},q_{2},q_{3}]\in\mathbb{Q}$. Then its magnitude is defined as
$|\tilde{q}|:=\sqrt{q_{0}^{2}+q_{1}^{2}+q_{2}^{2}+q_{3}^{2}}.$
And $\tilde{q}$ is invertible if and only if $|\tilde{q}|$ is positive. In this case, we have
$\tilde{q}^{-1}={\tilde{q}^{*}\over|\tilde{q}|^{2}}.$ (1)
A quaternion $\tilde{q}\in\mathbb{Q}$ is called a unit quaternion if $|\tilde{q}|=1$. Denote the set of all unit quaternions by $\mathbb{U}$. If $\tilde{p},\tilde{q}\in\mathbb{U}$, then $\tilde{p}\tilde{q}\in\mathbb{U}$. For any $\tilde{q}\in\mathbb{U}$, we have
$\tilde{q}\tilde{q}^{*}=\tilde{q}^{*}\tilde{q}=\tilde{1},$
i.e., $\tilde{q}$ is invertible and $\tilde{q}^{-1}=\tilde{q}^{*}$. A quaternion $\tilde{q}=[0,q_{1},q_{2},q_{3}]\in\mathbb{Q}$ is called a vector quaternion. A quaternion $\tilde{q}\in\mathbb{Q}$ is a vector quaternion if and only if $\tilde{q}=-\tilde{q}^{*}$. Suppose that $\tilde{q}$ is a vector quaternion and $\tilde{p}$ is a quaternion, then $\tilde{p}^{*}\tilde{q}\tilde{p}$ is still a vector quaternion. It is well-known that a unit quaternion can represent the rotation of a rigid body [6]. If a unit quaternion $\tilde{q}\in\mathbb{U}$ represents the rotation of a rigid body in the space, then we may write $\tilde{q}=[\cos\theta,\vec{l}\sin\theta]$, where $0\leq\theta<2\pi$ is the rotation angle of the rigid body, $\vec{l}$ is the rotation axis, which is a unit vector in $\mathbb{V}$. The logarithm of $\tilde{q}=[\cos\theta,\vec{l}\sin\theta]\in\mathbb{U}$ is a vector quaternion: $\ln\tilde{q}:=[0,\theta\vec{l}]$. A dual quaternion $\hat{q}=\tilde{q}+\tilde{q}_{d}\epsilon$ consists of two quaternions $\tilde{q}$, the standard part of $\hat{q}$, and $\tilde{q}_{d}$, the dual part of $\hat{q}$, with $\epsilon$ as the infinitesimal unit, satisfying $\epsilon\not=0$ but $\epsilon^{2}=0$. If both $\tilde{q}$ and $\tilde{q}_{d}$ are real numbers, then $\hat{q}$ is a dual number. We use a hat symbol to distinguish a dual quaternion. We denote the set of dual quaternions as $\hat{\mathbb{Q}}$. For $\hat{p}=\tilde{p}+\tilde{p}_{d}\epsilon\in\hat{\mathbb{Q}}$, if $\tilde{p}\not=0$, then we say that $\hat{p}$ is appreciable. Let $\hat{p}=\tilde{p}+\tilde{p}_{d}\epsilon,\hat{q}=\tilde{q}+\tilde{q}_{d}\epsilon\in\hat{\mathbb{Q}}$. Then the sum of $\hat{p}$ and $\hat{q}$ is
$\hat{p}+\hat{q}=(\tilde{p}+\tilde{q})+(\tilde{p}_{d}+\tilde{q}_{d})\epsilon$
and the product of $\hat{p}$ and $\hat{q}$ is
$\hat{p}\hat{q}=\tilde{p}\tilde{q}+(\tilde{p}\tilde{q}_{d}+\tilde{p}_{d}\tilde{q})\epsilon.$
Again, in general, $\hat{p}\hat{q}\not=\hat{q}\hat{p}$, but for any $\hat{p},\hat{q},\hat{s}\in\hat{\mathbb{Q}}$, we have
$(\hat{p}\hat{q})\hat{s}=\hat{p}(\hat{q}\hat{s}).$
The conjugate of $\hat{p}=\tilde{p}+\tilde{p}_{d}\epsilon$ is $\hat{p}^{*}=\tilde{p}^{*}+\tilde{p}_{d}^{*}\epsilon$. The magnitude of $\hat{p}$ is defined [11] as
$|\hat{p}|:=\left\{\begin{aligned} |\tilde{p}|+{(\tilde{p}\tilde{p}_{d}^{*}+\tilde{p}_{d}\tilde{p}^{*})\over 2|\tilde{p}|}\epsilon,&\ {\rm if}\ \tilde{p}\not={\tilde{0}},\\ |\tilde{p}_{d}|\epsilon,&\ {\rm otherwise},\end{aligned}\right.$ (2)
which is a dual number. In [11], a total order of dual numbers is defined. This is useful in the positive semidefiniteness theory of dual quaternion Hermitian matrices. Suppose that $\hat{p}=p+p_{d}\epsilon$ and $\hat{q}=q+p_{d}\epsilon$ are two dual numbers. Then $p,q,p_{d}$ and $q_{d}$ are real numbers. If $p>q$, or ${p}=q$ and $p_{d}>q_{d}$, then we say that $\hat{p}>\hat{q}$. In this way, we have positive, negative, nonpositive and nonnegative dual numbers. A dual vector $\hat{\vec{v}}=\vec{v}+\vec{v}_{d}\epsilon$ is a generalization of a dual number, whose standard part $\vec{v}$ and dual part $\vec{v}_{d}$ are both three-dimensional real vectors, i.e., $\vec{v},\vec{v}_{d}\in\mathbb{V}$. Then, a dual quaternion $\hat{q}$ can also be written as $\hat{q}=[\hat{s},\hat{\vec{v}}]$, where $\hat{s}$ is a dual number, and $\hat{\vec{v}}$ is a dual vector. A dual quaternion $\hat{p}=\tilde{p}+\tilde{p}_{d}\epsilon$ is called a unit dual quaternion if $\tilde{p}$ is a unit quaternion and
${\tilde{p}\tilde{p}_{d}^{*}+\tilde{p}_{d}\tilde{p}^{*}=\tilde{0}.}$ (3)
The set of unit dual quaternions is denoted as $\hat{\mathbb{U}}$. A dual quaternion $\hat{p}=\tilde{p}+\tilde{p}_{d}\epsilon$ is called a vector dual quaternion if both $\tilde{p}$ and $\tilde{p}_{d}$ are vector quaternions. A unit dual quaternion $\hat{q}=\tilde{q}+\tilde{q}_{d}\epsilon$ can represent the movement of a rigid body as
$\hat{q}=\tilde{q}+{\epsilon\over 2}\tilde{q}\tilde{p}^{b},$ (4)
where $\tilde{q}$ is a unit quaternion, representing the rotation of the rigid body, and $\tilde{p}^{b}=[0,\vec{p}\,^{b}]$ is a vector quaternion, with ${\vec{p}\,^{b}}$ representing the translation of the rigid body, and the superscript $b$ relating to the body frame, which is attached to the rigid body [15]. We may use $\tilde{q}$ to represent the configuration of the rigid body. Then $\tilde{q}$ represents the attitude of the rigid body, and $\tilde{p}^{b}$ represents the position of the rigid body. We may also write a unit dual quaternion as
$\hat{q}=\left[\cos\left(\hat{\theta\over 2}\right),\left(\sin\left(\hat{\theta\over 2}\right)\right)\hat{\vec{n}}\right],$ (5)
where dual vector $\hat{\vec{n}}=\vec{n}+\epsilon(\vec{c}\times\vec{n})$ is the screw axis, $\vec{c}$ is the vector from the original position to the rotational center, and $\hat{\theta}=\theta+\epsilon d$ is the dual angle of the screw [17]. With the unit dual quaternion $\hat{q}$ expressed by (4) and (5), the logarithm of $\hat{q}$ is defined as
${\ln\hat{q}=\left[\hat{0},{1\over 2}\hat{\theta}\hat{\vec{n}}\right].}$ (6)
We also denote the approximated logarithm of $\hat{q}$ [14] as
$\mathrm{lna}\,\hat{q}={1\over 2}(\tilde{\theta}+\epsilon\tilde{p}^{b})$ (7)
where $\tilde{\theta}=[0,\theta\vec{n}]$, $\tilde{p}^{b}=2\tilde{q}_{s}^{*}\tilde{q}_{d}$.
2.2 Dual Quaternion Vectors and Dual Quaternion Matrices
Denote $\hat{\mathbf{x}}=(\hat{x}_{1},\cdots,\hat{x}_{n})^{\top}\in{\hat{\mathbb{Q}}}^{n}$ for dual quaternion vectors. We say that $\hat{\mathbf{x}}\in\hat{\mathbb{Q}}^{n}$ is appreciable if at least one of its components is appreciable. We may also write
$\hat{\mathbf{x}}=\tilde{\mathbf{x}}+\tilde{\mathbf{x}}_{d}\epsilon,$
where $\tilde{\mathbf{x}},\tilde{\mathbf{x}}_{d}\in{\mathbb{Q}}^{n}$. Define ${\bar{\hat{\mathbf{x}}}}:=(\hat{x}_{1}^{*},\dots,\hat{x}_{n}^{*})^{\top}$ and $\hat{\mathbf{x}}^{*}:=(\hat{x}_{1}^{*},\dots,\hat{x}_{n}^{*})$. The $2$-norm of $\hat{\mathbf{x}}\in\hat{\mathbb{Q}}^{n}$ is defined as
$\|\hat{\mathbf{x}}\|_{2}=\left\{\begin{aligned} \sqrt{\sum_{i=1}^{n}|\hat{x}_{i}|^{2}},&\ {\rm if}\ \hat{\mathbf{x}}\not=\mathbf{0},\\ \|\hat{\mathbf{x}}_{d}\|_{2}\epsilon,&\ {\rm otherwise.}\end{aligned}\right.$ (8)
If $\|\hat{\mathbf{x}}\|_{2}=1$, then we say that $\hat{\mathbf{x}}$ is a unit dual quaternion vector. Then if $\hat{\mathbf{x}}\in\hat{\mathbb{U}}^{n}$, we have $\|\hat{\mathbf{x}}\|_{2}=\sqrt{n}$. If $\hat{\mathbf{x}}^{(1)},\dots,\hat{\mathbf{x}}^{(n)}\in\hat{\mathbb{Q}}^{n}$ and $\left(\hat{\mathbf{x}}^{(i)}\right)^{*}\hat{\mathbf{x}}^{(j)}=\delta_{ij}$ for $i,j=1,\cdots,n$, where $\delta_{ij}$ is the Kronecker symbol, then we say that $\{\hat{\mathbf{x}}^{(1)},\dots,\hat{\mathbf{x}}^{(n)}\}$ is an orthonormal basis of ^$\hat{\mathbb{Q}}^{n}$. By [12], we have the following proposition.
Proposition 2.1.
  • (i)
    $\|\hat{\mathbf{x}}\|_{2}\geq 0$, and $\|\hat{\mathbf{x}}\|_{2}=0$ if and only if $\hat{\mathbf{x}}=\hat{\mathbf{0}}$;
  • (ii)
    $\|\hat{q}\hat{\mathbf{x}}\|_{2}=|\hat{q}|\|\hat{\mathbf{x}}\|_{2}$;
  • (iii)
    $\|\hat{\mathbf{x}}+\hat{\mathbf{y}}\|_{2}\leq\|\hat{\mathbf{x}}\|_{2}+\|\hat{\mathbf{y}}\|_{2}$.
The collection of $m\times n$ dual quaternion matrices is denoted by ^$\hat{\mathbb{Q}}^{m\times n}$. A dual quaternion matrix $\hat{A}=(\hat{a}_{ij})\in\hat{\mathbb{Q}}^{m\times n}$ can be denoted as
$\hat{A}=\tilde{A}+\tilde{A}_{d}\epsilon,$ (9)
where $\tilde{A},\tilde{A}_{d}\in{\mathbb{Q}}^{m\times n}$. We call $\tilde{A}$ and $\tilde{A}_{d}$ the standard part and the infinitesimal part of $\hat{A}$, respectively. The transpose of $\hat{A}$ is $\hat{A}^{\top}=(\hat{a}_{ji})$. The conjugate transpose of $\hat{A}$ is $\hat{A}^{*}=(\hat{a}_{ji}^{*})$. Let $\hat{A}\in\hat{\mathbb{Q}}^{m\times n}$ and $\hat{B}\in\hat{\mathbb{Q}}^{n\times r}$. Then we have $(\hat{A}\hat{B})^{*}=\hat{B}^{*}\hat{A}^{*}$. Given a square dual quaternion matrix $\hat{A}\in\hat{\mathbb{Q}}^{n\times n}$, it is called invertible (nonsingular) if $\hat{A}\hat{B}=\hat{B}\hat{A}=I_{n}$ for some $\hat{B}\in\hat{\mathbb{Q}}^{n\times n}$, where $I_{n}$ is the $n\times n$ identity matrix. Such $\hat{B}$ is unique and denoted by $\hat{A}^{-1}$. Square dual quaternion matrix $\hat{A}$ is called Hermitian if $\hat{A}^{*}=\hat{A}$. Write ${\hat{A}}=\tilde{A}+\tilde{A}_{d}\epsilon\in\hat{\mathbb{Q}}^{n\times n}$. Then $\hat{A}$ is Hermitian if and only if both $\tilde{A}$ and $\tilde{A}_{d}$ are quaternion Hermitian matrices. A dual quaternion Hermitian matrix $\hat{A}\in{\hat{\mathbb{Q}}}^{n\times n}$ is called positive semidefinite if for any $\hat{\mathbf{x}}\in{\hat{\mathbb{Q}}}^{n}$, $\hat{\mathbf{x}}^{*}\hat{A}\hat{\mathbf{x}}\geq 0$; $\hat{A}$ is called positive definite if for any $\hat{\mathbf{x}}\in{\hat{\mathbb{Q}}}^{n}$ with $\hat{\mathbf{x}}$ being appreciable, we have $\hat{\mathbf{x}}^{*}\hat{A}\hat{\mathbf{x}}>0$ and is appreciable. Square dual quaternion matrix $\hat{A}$ is called unitary if $\hat{A}^{*}\hat{A}=I_{n}$. Apparently, $\hat{A}\in\hat{\mathbb{Q}}^{n\times n}$ is unitary if and only if its column vectors form an orthonormal basis of ^$\hat{\mathbb{Q}}^{n}$. Let $k\leq n$. We say that $\hat{A}\in\hat{\mathbb{Q}}^{n\times k}$ is partially unitary if its column vectors are unit vectors and orthogonal to each other. The following propositions and corollary are from [10].
Proposition 2.2.
Suppose that $\hat{A}\in\hat{\mathbb{Q}}^{n\times n}$ is unitary. Then $\hat{A}\hat{A}^{*}=I_{n}$.
Proposition 2.3.
Suppose that $\hat{U}\in\hat{\mathbb{Q}}^{n\times k}$ is partially unitary, $\hat{U}=\tilde{U}+\tilde{U}_{d}\epsilon$, where $\tilde{U},\tilde{U}_{d}\in{\mathbb{Q}}^{n\times k}$, $k\leq n$. Then $\tilde{U}$ is a quaternion partially unitary matrix, and
$\tilde{U}^{*}\tilde{U}_{d}+\tilde{U}_{d}^{*}\tilde{U}=O.$
Proposition 2.4.
Suppose that $\hat{U}\in\hat{\mathbb{Q}}^{n\times k}$ is partially unitary, where $k<n$. Then there is a vector $\hat{\mathbf{v}}\in\hat{\mathbb{Q}}^{n}$ such that $(\hat{U},\hat{\mathbf{v}})\in{\hat{\mathbb{Q}}}^{n\times(k+1)}$ is partially unitary.
Corollary 2.5.
Suppose that $\hat{U}\in\hat{\mathbb{Q}}^{n\times k}$ is partially unitary, $k<n$. Then there is $\hat{V}\in\hat{\mathbb{Q}}^{n\times(n-k)}$ such that $(\hat{U},\hat{V})\in\hat{\mathbb{Q}}^{n\times n}$ is unitary.
2.3 Eigenvalues and Singular Values of Dual Quaternion Matrices
The theory of eigenvalues of dual quaternion Hermitian matrices, and the theory of singular values of dual quaternion matrices were established in [12]. The contents of this subsection are from [12]. Suppose that $\hat{A}\in{\hat{\mathbb{Q}}}^{n\times n}$. If there are $\lambda\in\hat{\mathbb{Q}}$, $\hat{\mathbf{x}}\in{\hat{\mathbb{Q}}}^{n}$, where $\hat{\mathbf{x}}$ is appreciable, such that
$\hat{A}\hat{\mathbf{x}}=\hat{\mathbf{x}}\hat{\lambda},$ (10)
then we say that $\hat{\lambda}$ is a right eigenvalue of $\hat{A}$, with $\hat{\mathbf{x}}$ as an associated right eigenvector. If there are $\hat{\lambda}\in\hat{\mathbb{Q}}$, $\hat{\mathbf{x}}\in{\hat{\mathbb{Q}}}^{n}$, where $\hat{\mathbf{x}}$ is appreciable, such that
$\hat{A}\hat{\mathbf{x}}=\hat{\lambda}\hat{\mathbf{x}},$
then we say that $\hat{\lambda}$ is a left eigenvalue of $\hat{A}$, with $\hat{\mathbf{x}}$ as an associated left eigenvector. If $\hat{\lambda}$ is a dual number and a right eigenvalue of $\hat{A}$, then it is also a left eigenvalue of $\hat{A}$, as a dual number is commutative with a dual quaternion vector. In this case, we simply say that it is an eigenvalue of $\hat{A}$, with $\hat{\mathbf{x}}$ as an associated eigenvector.
Theorem 2.6.
Suppose that $\hat{A}=A+A_{d}\in{\hat{\mathbb{Q}}}^{n\times n}$ is a Hermitian matrix. Then there is a unitary matrix $\hat{U}\in{\hat{\mathbb{Q}}}^{n\times n}$ and a diagonal matrix $\hat{\Sigma}\in\hat{\mathbb{R}}^{n\times n}$ such that $\hat{\Sigma}=\hat{U}^{*}\hat{A}\hat{U}$, where
$\hat{\Sigma}\equiv{\rm diag}\left(\lambda_{1}+\lambda_{1,1}\epsilon,\cdots,\lambda_{1}+\lambda_{1,k_{1}}\epsilon,\lambda_{2}+\lambda_{2,1}\epsilon,\cdots,\lambda_{r}+\lambda_{r,k_{r}}\epsilon\right).$ (11)
with the diagonal entries of $\hat{\Sigma}$ being $n$ eigenvalues of $\hat{A}$,
$\hat{A}\hat{\mathbf{u}}_{i,j}=\hat{\mathbf{u}}_{i,j}(\lambda_{i}+\lambda_{i,j}\epsilon),$ (12)
for $j=1,\cdots,k_{i}$ and $i=1,\cdots,r$, $\hat{U}=(\hat{\mathbf{u}}_{1,1},\cdots,\hat{\mathbf{u}}_{1,k_{1}},\cdots,\hat{\mathbf{u}}_{r,k_{r}})$, $\lambda_{1}>\lambda_{2}>\cdots>\lambda_{r}$ are real numbers, $\lambda_{i}$ is a $k_{i}$-multiple right eigenvalue of $A$, $\lambda_{i,1}\geq\lambda_{i,2}\geq\cdots\geq\lambda_{i,k_{i}}$ are also real numbers, $\sum_{i=1}^{r}k_{i}=n$. Counting possible multiplicities $\lambda_{i,j}$, the form $\hat{\Sigma}$ is unique.
Theorem 2.7.
Suppose that $\hat{A}\in{\hat{\mathbb{Q}}}^{n\times n}$ is Hermitian. Then $\hat{A}$ has exactly $n$ eigenvalues, which are all dual numbers. There are also $n$ eigenvectors, associated with these $n$ eigenvalues. The Hermitian matrix $\hat{A}$ is positive semidefinite or definite if and only if all of these eigenvalues are nonnegative, or positive and appreciable, respectively.
Theorem 2.8.
Suppose that $\hat{B}\in{\hat{\mathbb{Q}}}^{m\times n}$. Then there exists a dual quaternion unitary matrix $\hat{V}\in{\hat{\mathbb{Q}}}^{m\times m}$ and a dual quaternion unitary matrix $\hat{U}\in{\hat{\mathbb{Q}}}^{n\times n}$ such that
$\hat{V}^{*}B\hat{U}=\left(\begin{aligned} \hat{\Sigma}_{t}\ \ \ &O\\ O\ \ \ &O\end{aligned}\right),$ (13)
where $\hat{\Sigma}_{t}\in\hat{\mathbb{R}}^{t\times t}$ is a diagonal matrix, taking the form
$\hat{\Sigma}_{t}={\rm diag}\left(\hat{\mu}_{1},\cdots,\hat{\mu}_{r},\cdots,\hat{\mu}_{t}\right),$
$r\leq t\leq\min\{m,n\}$, $\hat{\mu}_{1}\geq\hat{\mu}_{2}\geq\cdots\geq\hat{\mu}_{r}$ are positive appreciable dual numbers, and $\hat{\mu}_{r+1}\geq\cdots\geq\hat{\mu}_{t}$ are positive infinitesimal dual numbers. Counting possible multiplicities of the diagonal entries, the form $\hat{\Sigma}_{t}$ is unique.
We call $\hat{\mu}_{1},\cdots,\hat{\mu}_{t}$ and possibly $\hat{\mu}_{t+1}=\cdots=\hat{\mu}_{\min\{m,n\}}=0$, if $t<\min\{m,n\}$, the singular values of $\hat{B}$, $t$ the rank of $\hat{B}$, and $r$ the appreciable rank of $\hat{B}$.
3 Formation Control and Rank-One Unit Dual Quaternion Hermitian Matrices
Suppose that we have a group of $n$ rigid bodies in 3-D space. Then by [17, 15], they can be expressed by unit dual quaternions:
$\dot{\hat{q_{i}}}={1\over 2}\hat{q}_{i}{\hat{\xi}}_{i}^{b},$ (14)
for $i={1},\dots,n$, where the unit dual quaternion
$\hat{q}_{i}={\tilde{q}_{i}+{\epsilon\over 2}\tilde{q}_{i}\tilde{p}_{i}^{b}},$ (15)
represents the configuration of rigid body $i$ (including attitude $\tilde{q}_{i}$ and position $\tilde{p}_{i}^{b}$), and the dual vector quaternion
${\hat{\xi}}_{i}^{b}={\tilde{\omega}}_{i}^{b}+\epsilon(\dot{\tilde{p}}_{i}^{b}+{\tilde{\omega}}_{i}^{b}\times{\tilde{p}}_{i}^{b})$ (16)
is the twist of the rigid body $i$ (including the angular velocity ${\tilde{\omega}}_{i}^{b}$ and the linear velocity $\dot{\tilde{p}}_{i}^{b}$), the superscript $b$ relates to the body frame, which is attached to the rigid body [15]. Note that the limit of dual quaternion means that both the standard part and the dual part of the dual quaternion converge to the standard and dual part of this limit, respectively. This group of $n$ rigid bodies with the kinematics expressed by (14-16) is called the overall system. If for any pair of rigid bodies $i$ and $j$ in the overall system, the actual relative configuration $\hat{q}_{ij}$ converges to the desired relative configuration $\hat{q}_{d_{ij}}$ asymptotically, i.e., $\hat{q}_{ij}(t)\to\hat{q}_{d_{ij}}$, then we say that the overall system is pairwise asymptotically stable. Denote $\hat{Q}_{d}=(\hat{q}_{d_{ij}})$. Then this is an $n\times n$ dual quaternion matrix. Similarly, we may denote $\hat{Q}(t)=(\hat{q}_{ij}(t))$. Then it is also an $n\times n$ dual quaternion matrix. If
$\lim_{t\to\infty}\hat{Q}(t)=\hat{Q}_{d},$ (17)
then we have pairwise asymptotical stability of the overall system. Then we have the following theorem.
Theorem 3.1.
In the above scheme, $\hat{Q}(t)$ is a rank-one dual quaternion Hermitian matrix, with a unique nonzero eigenvalue $n$. All of its diagonal entries are $1$. Write its first row vector as
$\hat{\mathbf{q}}_{1}(t)^{\top}=(1,\hat{q}_{12}(t),\dots,\hat{q}_{1n}(t)).$ (18)
Then we have
$\hat{Q}(t)=\bar{\hat{\mathbf{q}}}_{1}(t)\hat{\mathbf{q}}_{1}(t)^{\top}.$ (19)
Then, if pairwise asymptotical stability of the overall system is required, $\hat{Q}_{d}$ must also be a rank-one dual quaternion Hermitian matrix, with a unique nonzero eigenvalue $n$. All of its diagonal entries are $1$. Write its first row vector as
$\hat{\mathbf{q}}_{d_{1}}^{\top}=(1,\hat{q}_{d_{12}},\dots,\hat{q}_{d_{1n}}).$ (20)
Then we have
$\hat{Q}_{d}=\bar{\hat{\mathbf{q}}}_{d_{1}}\hat{\mathbf{q}}_{d_{1}}^{\top}.$ (21)
Proof.
We know that
$\hat{q}_{ij}(t)=\hat{q}_{i}^{*}(t)\hat{q}_{j}(t),$
where $\hat{q}_{i}(t)$ is the configuration of rigid body $i$ at time $t$, for $i,j=1,\dots,n$. Then for $i=1,\dots,n$,
$\hat{q}_{ii}(t)=\hat{q}_{i}^{*}(t)\hat{q}_{i}(t)=1.$
This shows that all diagonal entries of $\hat{Q}(t)$ are $1$. Then we may write the first row vector of $\hat{Q}(t)$ as (18). For $i,j=1,\dots,n$, we have
$\hat{q}_{ij}(t)=\hat{q}_{i}^{*}(t)\hat{q}_{j}(t)=\hat{q}_{1i}^{*}(t)\hat{q}_{1j}(t).$ (22)
Here, the second equality follows from $\hat{q}_{1i}(t)=\hat{q}_{1}^{*}(t)\hat{q}_{i}(t)$ for all $i=1,\dots,n$. This implies (19). By (19), Corollary 2.5, Theorems 2.6 and 2.8, $\hat{Q}(t)$ is a rank-one unit dual quaternion Hermitian matrix, with a unique nonzero eigenvalue $n$. By (17), if pairwise asymptotical stability of the overall system is required, we have the remaining conclusion. ∎
This theorem shows that $\hat{Q}_{d}$ cannot be arbitrarily given, if pairwise asymptotical stability of the overall system is targeted. We now reconsider this overall system of $n$ rigid bodies. If for $j=2,\dots,n$, the actual relative configuration $\hat{q}_{1j}$ converges to the desired relative configuration $\hat{q}_{d_{1j}}$ asymptotically, i.e., $\hat{q}_{1j}(t)\to\hat{q}_{d_{1j}}$, then we say that the overall system is row-one asymptotically stable.
Theorem 3.2.
The overall system is pairwise asymptotically stable if and only if it is row-one asymptotically stable.
Proof.
As the conditions of row-one asymptotical stability are only a proper subset of the conditions of pairwise asymptotical stability, if we have pairwise asymptotical stability, then we have row-one asymptotical stability. On the other hand, assume row-one asymptotical stability is reached. For $i=2,\dots,n$ and $j=1,\dots,n$, let
$\hat{q}_{d_{ij}}=\hat{q}_{d_{1i}}^{*}\hat{q}_{d_{1j}}.$
Then row-one asymptotical stability implies pairwise asymptotical stability by (22).
This theorem reduces pairwise asymptotical stability, which has $n^{2}$ conditions, to row-one asymptotical stability, which has only $n-1$ conditions. In [15], a control strategy via a rooted directed tree was used. We may connect the overall system by a rooted directed tree $T$, with the $n$ rigid bodies as its nodes. Without loss of generality, we may choose rigid body $1$ as the root. Then $T$ has $n-1$ directed arcs $(i,j)$, where $i$ is the tail of the directed arc, and $j$ is the head of the directed arc. Then the root $1$ is the only node that is not a head, while a node, which is not a tail, is a leaf of the tree $T$. If for all $n-1$ directed arcs $(i,j)$ in $T$, the actual relative configuration $\hat{q}_{ij}$ converges to the desired relative configuration $\hat{q}_{d_{ij}}$ asymptotically, i.e., $\hat{q}_{ij}(t)\to\hat{q}_{d_{ij}}$, then we say that the overall system is tree-wise asymptotically stable. We may see that row-one asymptotical stability is a special tree-wise asymptotical stability. Similar to Theorem 3.2, we have the following theorem.
Theorem 3.3.
The overall system is pairwise asymptotically stable if and only if it is tree-one asymptotically stable for a certain rooted direct tree $T$ covering the overall system.
This theorem generalizes Theorem 3.2.
4 Dual Quaternion Laplacian
Consider the formation control problem of $n$ rigid bodies. Then these $n$ rigid bodies can be described by a graph $G=(V,E)$ with $n$ vertices and $m$ edges. For two rigid bodies $i$ and $j$ in $V$, if rigid body $i$ can sense rigid body $j$, then edge $(i,j)\in E$. As studied in [13], we assume that any pair of these rigid bodies are mutual visual, i.e., rigid body $i$ can sense rigid body $j$ if and only if rigid body $j$ can sense rigid body $i$. Furthermore, we assume that $G$ is connected in the sense that for any node pair $i$ and $j$, either $(i,j)\in E$, or there is a path connecting $i$ and $j$ in $G$, i.e., there are nodes $i_{i},\dots,i_{k}\in V$ such that $(i,i_{1}),\dots,(i_{k},j)\in E$. Suppose that for each $(i,j)\in E$, we have a desired relative configuration from rigid body $i$ to $j$ as $\hat{q}_{d_{ij}}\in\hat{\mathbb{U}}$. We say that the desired relative configurations $\left\{\hat{q}_{d_{ij}}:(i,j)\in E\right\}$ is reasonable if and only if there is a desired formation $\hat{\mathbf{q}}_{d}\in\hat{\mathbb{U}}^{n\times 1}$, which satisfies
$\hat{q}_{d_{ij}}=\hat{q}_{d_{i}}^{*}\hat{q}_{d_{j}},$ (23)
for all $(i,j)\in E$.
Theorem 4.1.
Suppose that we have an undirected graph $G=(V,E)$, which is bidirectional and connected. The desired relative configurations $\left\{\hat{q}_{d_{ij}}:(i,j)\in E\right\}$ is reasonable if and only if
  • for all $(i,j)\in E$, we have
    $\hat{q}_{d_{ji}}=\hat{q}_{d_{ij}}^{*};$ (24)
  • for any cycle $\left\{j_{1},\dots,j_{k}\right\}$, with $j_{k+1}=j_{1}$, of $G$, we have
    $\prod_{i=1}^{k}\hat{q}_{d_{j_{i}j_{i+1}}}=1.$ (25)
Furthermore, let ${\hat{\mathbf{q}}}_{0}$ be a desired formation satisfying (23). Then the set of all the desired formations satisfying (23) is
$\left\{\hat{c}\hat{\mathbf{q}}_{0}:\hat{c}\in\hat{\mathbb{U}}\right\}.$ (26)
Proof.
Suppose that the desired relative configurations $\left\{\hat{q}_{d_{ij}}:(i,j)\in E\right\}$ is reasonable. Then there is a desired formation $\hat{\mathbf{q}}_{d}\in\hat{\mathbb{U}}^{n\times 1}$, which satisfies (23) for all $(i,j)\in E$. Then for all $(i,j)\in E$, we have $\hat{q}_{d_{ji}}=\hat{q}_{d_{j}}^{*}\hat{q}_{d_{i}}=\hat{q}_{d_{ij}}^{*}$, which implies (24). On the cycle $\left\{j_{1},\dots,j_{k}\right\}$, we have $\hat{q}_{d_{j_{i+1}}}=\hat{q}_{d_{j_{i}}}\hat{q}_{d_{{j_{i}}{j_{i+1}}}}$ for $i=1,\dots,k$. This implies (25). On the other hand, suppose that the desired relative configurations $\left\{\hat{q}_{d_{ij}}\right\}$ satisfy (24) for all $(i,j)\in E$, and (25) for all cycles in $G$. We wish to show that there is a desired formation $\hat{\mathbf{q}}_{d}\in\hat{\mathbb{U}}^{n\times 1}$, which satisfies (23) for all $(i,j)\in E$. We count two edges $(i,j)$ and $(j,i)$ as one bidirectional edge pair. Thus, $G$ has $m_{0}={m\over 2}$ bidirectional edge pairs. By graph theory, since $G$ is bidirectional and connected, we have $m_{0}\geq n-1$. Let $M=m_{0}-n+1$. We now show by induction on $M$, that there is a desired formation $\hat{\mathbf{q}}_{d}\in\hat{\mathbb{U}}^{n\times 1}$, which satisfies (23) for all $(i,j)\in E$. (i) We first show this for $M=0,i.e.,m_{0}=n-1$. Then $G$ is a bidirectional tree. Without loss of generality, we pick node $1$ as the root of this tree. Let $\hat{q}_{d_{1}}=\hat{1}$. Each node in $\{2,\dots,n\}$ has one unique father node. Denote $V_{(1)}$ as the set of children nodes of the root. Let $\hat{q}_{d_{i}}=\hat{q}_{d_{1}}\hat{q}_{d_{1i}}=\hat{q}_{d_{1i}}$ for all $i\in V_{(1)}$. Similarly, denote $V_{(j)}$ as the set of children nodes of $V_{{(j-1)}}$. Then $\hat{q}_{d_{i}}=\hat{q}_{d_{f_{i}}}\hat{q}_{d_{f_{i}i}}$ for all $i\in V_{(j)}$, where $f_{i}$ is the father node of node $i$. We repeat this process until $V=\cup_{j}V_{(j)}$. By (24), equation (23) is satisfied for all $(i,j)\in E$. (ii) We now assume that this claim is true for $M=M_{0}$, and prove that it is true for $M=M_{0}+1$. Then $M>m_{0}-n+1$. This implies that there is at least one cycle $\left\{j_{1},\dots,j_{k}\right\}$, with $j_{k+1}=j_{1}$, of $G$, such that the $k$ nodes $j_{1},\dots,j_{k}$ are all distinct and $k\geq 3$. Delete the bidirectional edge pair $(j_{1},j_{k})$ and $(j_{k},j_{1})$. We get an undirectional graph $G_{0}=(V,E_{0})$. Then $G_{0}$ is still bidirectional and connected, and the desired relative configurations $\left\{\hat{q}_{d_{ij}}:(i,j)\in E_{0}\right\}$ satisfy (24) for all $(i,j)\in E_{0}$, and (25) for all cycles in $G_{0}$. By our induction assumption, there is a desired formation $\hat{\mathbf{q}}_{d}\in\hat{\mathbb{U}}^{n\times 1}$ such that (23) holds for all $(i,j)\in E_{0}$. Now, by (24) for $i=j_{1}$ and $j=j_{k}$, and (25) for this cycle $\left\{j_{1},\dots,j_{k}\right\}$, (23) also holds for $(j_{1},j_{k})$ and $(j_{k},j_{1})$. This proves this claim for $M=M_{0}+1$ in this case. Suppose that ${\hat{\mathbf{q}}}_{0}$ is a desired formation satisfying (23), and $\hat{c}\in\hat{\mathbb{U}}$. Then $\hat{c}{\hat{\mathbf{q}}}_{0}$ also satisfies (23). We may prove the other side of the last claim by induction as above. The proof is complete. ∎
Actually, we may regard (24) as a special case of (25) with $k=2$. The set (26) means that while the $n$ rigid bodies converge asymptotically to the desired relative formation, they may still have rotations and translations as a cohesive whole as in [8], or they have a lead which moves following some trajectories and the other rigid bodies follow that lead as in [15]. Now, the Laplacian matrix of $G$ is defined as follows,
$\hat{L}={D-\hat{A}},$ (27)
where $D$ is a diagonal real matrix whose $i$-th diagonal element is equal to the degree $d_{i}$ of the $i$-th vertex, and ${\hat{A}}=(\hat{a}_{ij})$,
$\hat{a}_{ij}=\left\{\begin{array}[]{cl}{\hat{q}_{d_{ij}}},&\text{if }(i,j)\in E,\\ 0,&\text{otherwise},\end{array}\right.$
where $\hat{q}_{d_{ij}}$ is the desired relative configuration on the edge $(i,j)\in E$. Now for $(i,j)\in E$, let $\hat{a}_{ij}\in\hat{\mathbb{U}}$ be the desired relative configuration of the rigid bodies $i$ and $j$. Then $\hat{A}$ is the desired relative configuration adjacency matrix. By (24), both $\hat{A}$ and $\hat{L}$ are dual quaternion Hermitian matrices. This is important for their spectral analysis, as by [12], an $n\times n$ dual quaternion Hermitian matrix has exactly $n$ eigenvalues, which are dual numbers. On the other hand, if $\hat{A}$ and $\hat{L}$ are not Hermitian matrices, then their spectral analysis may be very complicated. Then we have
$\hat{L}\bar{\hat{\mathbf{q}}}_{d}=\hat{\mathbf{0}},$ (28)
where $\bar{\hat{\mathbf{q}}}_{d}=[{\hat{q}}^{*}_{d_{1}},\cdots,{\hat{q}}^{*}_{d_{n}}]^{\top}$ is the conjugate of $\hat{\mathbf{q}}_{d}$. With the above construction, we have the following theorem.
Theorem 4.2.
Suppose that we have a bidirectional and connected graph $G=(V,E)$ with reasonable desired relative configurations $\left\{\hat{q}_{d_{ij}}:(i,j)\in E\right\}$ on $G$. Let the dual quaternion Laplacian $\hat{L}$ be constructed as above. Then a formation ${\hat{\mathbf{q}}}_{d}$ is a desired formation satisfying (23), if and only if it is in $\hat{\mathbb{U}}^{n\times 1}$ and satisfies (28). Furthermore, let $\hat{Q}=\mathrm{diag}(\hat{\mathbf{q}}_{d})$. Then $\hat{Q}$ is a unitary matrix, and
$\hat{Q}{\hat{L}}\hat{Q}^{*}=L=D-A,$ (29)
where $D$, $A$, and $L$ are the degree matrix, the unweighted adjacency matrix, and the unweighted Laplacian matrix of $G$. Therefore, all the eigenvalues of $\hat{L}$ are real numbers. One of them is zero. The others are positive. The conjugates of all the desired formation vectors are the eigenvectors of $\hat{L}$, corresponding to the zero eigenvalue of $\hat{L}$.
Proof.
Suppose that ${\hat{\mathbf{q}}}_{d}$ is a desired formation. Then it satisfies (23). Then the left of the $i$th equation of (28) satisfies
$d_{i}\hat{q}_{d_{i}}^{*}-\sum_{(i,j)\in E}\hat{q}_{d_{ij}}\hat{q}_{d_{j}}^{*}=d_{i}\hat{q}_{d_{i}}^{*}-\sum_{(i,j)\in E}\hat{q}_{d_{i}}^{*}\hat{q}_{d_{j}}\hat{q}_{d_{j}}^{*}=0.$
Thus, it satisfies (28). From this, we have (29). Since $G$ is connected, by spectral graph theory, the eigenvalues of $L$ are all positive numbers, except one zero eigenvalue. But (29) implies that $\hat{L}$ and $L$ are similar, by [12], the eigenvalues of $\hat{L}$ are the same as the eigenvalues of $L$. Thus, they are all positive numbers, except one zero eigenvalue. On the other hand, suppose that $\hat{\mathbf{q}}\in\hat{\mathbb{U}}^{n\times 1}$ and satisfies (28). We now need to show that it satisfies (23). Then, by (29), we have ${\hat{L}}=\hat{Q}^{*}L\hat{Q}$ and
$\hat{L}\bar{\hat{\mathbf{q}}}=\hat{\mathbf{0}}\Leftrightarrow\hat{Q}^{*}L\hat{Q}\bar{\hat{\mathbf{q}}}=\hat{\mathbf{0}}\Leftrightarrow L\hat{Q}\bar{\hat{\mathbf{q}}}=\hat{\mathbf{0}}$
By spectral graph theory, the zero eigenvalue of $L$ has an eigenvector $\hat{\bf 1}$. Therefore,
$\hat{Q}\bar{\hat{\mathbf{q}}}={\hat{\bf 1}}\hat{c}$
for any $c\in\hat{\mathbb{Q}}$. Hence,
$\bar{\hat{\mathbf{q}}}=\hat{Q}^{*}{\hat{\bf 1}}\hat{c}=\bar{\hat{\mathbf{q}}}_{d}\hat{c}.$
Therefore, $\hat{\mathbf{q}}=\hat{c}^{*}\hat{\mathbf{q}}_{d}$ and $\hat{c}\in\hat{\mathbb{U}}$. The sufficient condition holds from Theorem 4.1. This also shows that the conjugates of all the desired formation vectors are the eigenvectors of $\hat{L}$, corresponding to the zero eigenvalue of $\hat{L}$. The proof is completed. ∎
This theorem establishes the relation between desired formations and the null space of the dual quaternion Laplacian, and the relation between the dual quaternion Laplacian and the ordinary Laplacian of $G$.
5 Control Law
One possible way is to follow [8]. Suppose that in the above formation control problem each rigid body $i$ has a point kinematic model given by the single integrator
$\dot{\hat{z}}_{i}=\hat{u}_{i},\ i=1,\cdots,n,$ (30)
where $\hat{\mathbf{z}}=\hat{\mathbf{z}}(t)\in\hat{\mathbb{U}}^{n\times 1}$ is the unit dual quaternion vector of the observed configurations of these $n$ rigid bodies at time $t$. Consider the following control law
$\hat{u}_{i}=\hat{k}_{i}\left(\sum_{j\in\mathcal{N}_{i}}{\hat{q}_{d_{ij}}}\hat{z}_{j}-d_{i}\hat{z}_{i}\right)$ (31)
or equivalently, $\hat{\mathbf{u}}=-\hat{K}\hat{L}\hat{\mathbf{z}}$. Here, $\mathcal{N}_{i}$ is the set of neighbors of $i$, $d_{i}$ is the degree of $i$, and $\hat{K}=diag(\hat{k}_{1},\cdots,\hat{k}_{n})$ is the stabilizing matrix applied to control the eigenvalues of $\hat{K}\hat{L}$. By [8], if the eigenvalues of the closed-loop system are at the open left half complex plane, the system will have asymptotic stability. However, the eigenvalues of the closed-loop system are required to be far away from the imaginary axis, if in addition robustness is required. The overall closed-loop dynamics of $n$ agents becomes
$\dot{\hat{\mathbf{z}}}=-\hat{K}\hat{L}\hat{\mathbf{z}}.$ (32)
Theorem 5.1.
Suppose that we have a bidirectional and connected graph $G=(V,E)$ with reasonable desired relative configurations $\left\{\hat{q}_{d_{ij}}:(i,j)\in E\right\}$ on $G$. Let $\hat{\mathbf{q}}_{d}\in\hat{\mathbb{U}}^{n\times 1}$ be a desired formation satisfying (23). If there exists a positive diagonal matrix $K\in\mathbb{R}^{n\times n}$ such that $\hat{K}=\hat{Q}^{*}K\hat{Q}$, where $\hat{Q}=diag(\hat{\mathbf{q}}_{d})$, then the observed configuration vector $\hat{\mathbf{z}}=\hat{\mathbf{z}}(t)$ under control law (31) globally asymptotically converges to $\bar{\hat{\mathbf{q}}}_{d}\hat{c}$ with $-^\hat{c}\in{\hat{\mathbb{Q}}}$.
Proof.
From the assumptions we have
$\hat{K}\hat{L}=(\hat{Q}^{*}K\hat{Q})(\hat{Q}^{*}L\hat{Q})=\hat{Q}^{*}KL\hat{Q}.$
Since $K$ is a positive definite diagonal matrix and $KL=K^{\frac{1}{2}}(K^{\frac{1}{2}}LK^{\frac{1}{2}})K^{-\frac{1}{2}}$, all eigenvalues of $KL$ are real numbers, and all eigenvalues of $KL$ are positive except one zero eigenvalue. Denote the eigenvalue decomposition of $K^{\frac{1}{2}}LK^{\frac{1}{2}}$ as $U\Sigma U^{\top}$. Without loss of generalization, let $\Sigma=diag(0,\Lambda)$. Then $\Lambda$ is a diagonal matrix and all diagonal elements are positive. Let $V=cK^{\frac{1}{2}}U$. Then there is
$V^{-1}KLV=\begin{bmatrix}0&0\\ 0&\Lambda\end{bmatrix}.$ (33)
Here, $c=\frac{1}{{\bf 1}^{\top}K^{\frac{1}{2}}u_{1}}$, $u_{1}$ is the first column of $U$. Therefore, the first column of $V\in\mathbb{R}^{n\times n}$ is $\mathbf{1}_{n}$ and all eigenvalues of $\Lambda$ are positive. For any initial configuration $\hat{\mathbf{z}}(0)\in\hat{\mathbb{U}}^{n}$, define
${\hat{\mathbf{y}}}=\begin{bmatrix}\hat{y}_{1}\\ {\hat{\mathbf{y}}_{2}}\end{bmatrix}=V^{-1}\hat{Q}\hat{\mathbf{z}}.$ (34)
Then there is $\hat{\mathbf{y}}(0)=V^{-1}\hat{Q}\hat{\mathbf{z}}(0)\in\hat{\mathbb{Q}}^{n}$ and
$\begin{bmatrix}\dot{\hat{y}}_{1}\\ \dot{{\hat{\mathbf{y}}}}_{2}\end{bmatrix}=-\begin{bmatrix}0&0\\ 0&\Lambda\end{bmatrix}\begin{bmatrix}\hat{y}_{1}\\ {\hat{\mathbf{y}}_{2}}\end{bmatrix}.$ (35)
It follows that $\hat{y}_{1}(t)=\hat{y}_{1}(0)$ and $\hat{\mathbf{y}}_{2}(t)=\exp(-\Lambda t)\hat{\mathbf{y}}_{2}(0)$ asymptotically converges to a zero vector as $t\rightarrow\infty$. Therefore, $\hat{\mathbf{z}}(t)={\hat{Q}^{*}V\hat{\mathbf{y}}(t)}$ globally asymptotically converges to $\bar{\hat{\mathbf{q}}}_{d}\hat{y}_{1}(0)$. This completes the proof.
Denote the limit of $\hat{\mathbf{z}}(t)$ as $\hat{\mathbf{z}}^{\infty}=\bar{\hat{\mathbf{q}}}_{d}\hat{y}_{1}(0)$. For an arbitrary initial configuration $\hat{\mathbf{z}}(0)\in\hat{\mathbb{U}}^{n}$, it is possible that $\hat{\mathbf{z}}^{\infty}\notin\hat{\mathbb{U}}^{n}$ , hence is not a configuration. However, if $\hat{y}_{1}(0)=(V^{-1})_{1}\hat{Q}\hat{\mathbf{z}}(0)$ is an invertible dual quaternion number, or equivalently, $(V^{-1})_{1}\tilde{Q}_{s}\tilde{z}_{s}(0)$ is an invertible quaternion number, we can define the final configuration as
$\hat{\mathbf{z}}^{c}=\hat{\mathbf{z}}^{\infty}(\hat{z}_{1}^{\infty})^{-1}\in\hat{\mathbb{U}}^{n}.$
We may verify that $\hat{\mathbf{z}}^{c}$ belongs to the set of all the desired formations satisfying (23) defined by (26).
Remark 5.2.
In the dynamics system defined by (32), it is possible that $\hat{\mathbf{z}}(t)\notin\hat{\mathbb{U}}^{n}$ even if $\hat{\mathbf{z}}(0)\in\hat{\mathbb{U}}^{n}$. In practice, we can project $\hat{\mathbf{z}}(t)$ to a point in $\hat{\mathbb{U}}^{n}$. Therefore, defining a dynamics system such that $\hat{\mathbf{z}}(t)\in\hat{\mathbb{U}}^{n}$ for all $t$ is an important topic.
The stabilizing matrix $\hat{K}\in\hat{\mathbb{Q}}^{n\times n}$ is a positive diagonal dual quaternion matrix satisfying $\hat{K}=\hat{Q}^{*}K\hat{Q}$, where $K\in\mathbb{R}^{n\times n}$ is a positive diagonal real matrix. Consider two possible tasks. The first task is to find $K$ such that the second smallest eigenvalues of $KL$ is larger than a tolerance $\delta$. Since
$k_{\min}\lambda_{i}(L)\leq\lambda_{i}(KL)\leq k_{\max}\lambda_{i}(L),$
where $k_{\min}=\min_{i}k_{i}$ and $k_{\max}=\max_{i}k_{i}$, we can set $k_{\min}=\frac{\lambda_{n-1}(L)}{\delta}$, where $\lambda_{n-1}(L)$ is the second smallest eigenvalues of $L$. It follows from spectral graph theory, $\lambda_{n-1}(L)$ is the algebra connectivity of that graph. We can also add more edges to $G$ and increase the connectivity of the graph. The second task is to compute $K$ such that the eigenvalues of $KL$ are equal to the prescribed numbers $\{0,\sigma_{1},\dots,\sigma_{n-1}\}$. In this case, we can use numerical approaches. Specifically, we compute $k_{1},\cdots,k_{n}$ by solving the following system
$\det(\sigma_{i}I{-}KL)=0,\quad i=1,\cdots,n.$
Another way is to follow the tree approach of [15]. According to (25) of [15], a general formation control law for each rigid body $j$ in the rooted tree except the lead is as follows:
$\hat{\xi}_{j}^{b}=\left(\hat{q}_{d_{ij}}^{*}\hat{q}_{ij}\right)\left(\hat{\xi}_{i}^{b}+2\hat{k}_{j}{\mathrm{lna}}\left(\hat{q}_{d_{ij}}^{*}\hat{q}_{ij}\right)\right)\left(\hat{q}_{ij}^{*}\hat{q}_{d_{ij}}\right),$ (36)
where rigid body $i$ is the father node of rigid body $j$, $q_{ij}$ is the current relative configuration from rigid $i$ to $j$, and ${\mathrm{lna}}$ is the approximated logarithm defined in (7). Here, $\hat{q}_{d_{ij}}^{*}\hat{q}_{ij}$ is need, where $(i,j)$ is an arc on the control interconnection rooted tree.
6 Directed Graph
Consider the formation control problem of $n$ rigid bodies, which can be described by a directed (digraph) $G=(V,E)$ with $n$ nodes and $m$ arcs. For two rigid bodies $i$ and $j$ in $V$, if rigid body $i$ can sense rigid body $j$, then arc $(i,j)\in E$, where $i$ and $j$ are the tail and head of the arc $(i,j)$ respectively. It is possible that $(i,j)\in E$ but $(j,i)\not\in E$, i.e., rigid body $i$ can sense rigid body $j$ but rigid body $j$ cannot sense rigid body $i$. A digraph $G$ is strongly connected if for every ordered pair of vertices $i$ and $j$, there is a directed path from $i$ to $j.$ A digraph $G$ is unilaterally connected if for every ordered pair of vertices $i$ and $j$, there is a directed path from $i$ to $j$ or a directed path from $j$ to $i$. A digraph $G$ is weakly connected if the underlying undirected graph is connected. A digraph $G$ is not connected if it is not weakly connected. In general, a rooted tree is weakly connected by not strongly connected. In a rooted tree, if from the root to any other node of the tree there is a directed path, then this rooted tree is a directed rooted tree. The Laplacian matrix of a directed graph $G$ is still defined by (27). However, for a directed graph, neither $\hat{A}$ nor $\hat{L}$ is a dual quaternion Hermitian matrix. In this case, their spectral analysis is complicated. The spectral theory of the Laplacian matrix of a directed graph $G$ is important. Let $L$ be the unweighted Laplacian of $G$. Some basic properties of $L$ are as follows.
  • 00 is an eigenvalue of $L$ and the associated eigenvector is $\bf 1$.
  • The number of zero eigenvalues of $L$ is equal to the in-forest dimension [2] or the number of reaches [1] of $G$. An in-forest is a spanning converging forest. An in-forest $F$ of a digraph $G$ is called a maximum in-forest of $G$ if $G$ has no in-forest with a greater number of arcs than in $F$. The number of maximum in-forests is called the in-forest dimension of $G$ and denoted by $d$. The number of arcs in any maximum in-forest is obviously $n-d$. In Proposition 12(ii) of [2], it was shown that rank${(L)}=n-d$. In [1], they write $j\rightsquigarrow i$ if there exists a directed path from vertex $j$ to vertex $i$. For any vertex $j$, define the reachable set $R(j)$ to be the set containing $j$ and all vertices $i$ such that $j\rightsquigarrow i.$ A maximal reachable set will be called a reach. In Theorem 3.2 of [1], it was shown that the algebraic and geometric multiplicity of 0 as an eigenvalue of $L$ equals the number of reaches of $G.$
  • All eigenvalues of $D^{-1}L$ are located in a disk of radius 1 in the complex plane centered at $01+j$ 0. This can be shown by applying Gershgorin’s theorem to the rows of $D^{-1}L$. Similarly, let $d_{\max}$ be the maximal degree of $G$. By Gershgorin’s theorem, all eigenvalues of $L$ are located in a disk of radius $d_{\max}$ in the complex plane centered at $d_{\max}+j0$.
Based on Proposition 12(ii) of [2] and Theorem 3.2 of [1], we present the following proposition and omit its proof here.
Proposition 6.1.
Suppose that we have a directed and connected graph $G=(V,E)$. If the in-forest dimension of $G$ or the number of reaches of $G$ is one, then $L$ has only one zero eigenvalue and its corresponding eigenvector is $\bf 1$.
We can verify that the in-forest dimension of a rooted tree and a strongly connected directed graph is one. Hence, a rooted tree and a strongly connected directed graph have only one right zero eigenvalue with right eigenvector $\bf 1$. This property is important for our theoretical analysis. Similar to Theorem 4.2, we have the following theorem.
Theorem 6.2.
Suppose that we have a directed and connected graph $G=(V,E)$ with reasonable desired relative configurations $\left\{\hat{q}_{d_{ij}}:(i,j)\in E\right\}$ on $G$. Assume that the in-forest dimension of $G$ is one. Let the dual quaternion Laplacian $\hat{L}$ be constructed as above. Then a formation ${\hat{\mathbf{q}}}_{d}$ is a desired formation satisfying (23), if and only if it is in $\hat{\mathbb{U}}^{n\times 1}$ and satisfies (28). Furthermore, let $\hat{Q}=\mathrm{diag}(\hat{\mathbf{q}}_{d})$. Then $\hat{Q}$ is a unitary matrix, and
$\hat{Q}{\hat{L}}\hat{Q}^{*}=L=D-A,$ (37)
where $D$, $A$, and $L$ are the degree matrix, the unweighted adjacency matrix, and the unweighted Laplacian matrix of $G$. Furthermore, all eigenvalues of $L$ are right eigenvalues of $\hat{L}$, which are complex numbers. Then zero is an eigenvalue of $\hat{L}$. The conjugates of all the desired formation vectors are eigenvectors of $\hat{L}$, corresponding to the zero eigenvalue of $\hat{L}$.
Proof.
The proof of the first claim and (37) is similar to the proof of Theorem 4.2. We do not go to the details here. Suppose that $\hat{\lambda}$ is an eigenvalue of $\hat{L}$ with an eigenvector $\hat{\mathbf{x}}$. From the definition, there is
$\hat{L}\hat{\mathbf{x}}=\hat{\mathbf{x}}\hat{\lambda}\Leftrightarrow\hat{Q}^{*}L\hat{Q}\hat{\mathbf{x}}=\hat{\mathbf{x}}\hat{\lambda}\Leftrightarrow L\hat{Q}\hat{\mathbf{x}}=\hat{Q}\hat{\mathbf{x}}\hat{\lambda}\Leftrightarrow L\hat{\mathbf{y}}=\hat{\mathbf{y}}\hat{\lambda},$
where $\hat{\mathbf{y}}=\hat{Q}\hat{\mathbf{x}}$. In other words, all eigenvalues of $L$ are also right eigenvalues of $\hat{L}$. By Lemma 6.1, the Laplacian matrix $L$ has only one zero eigenvalue and its corresponding eigenvector is $\bf 1$. Therefore, zero is also a right eigenvalue of $\hat{L}$ and its corresponding eigenvector is $\hat{Q}^{*}{\bf 1}=\bar{\hat{\mathbf{q}}}_{d}$. By Section 2, a real right eigenvalue of a dual quaternion matrix can be simply called an eigenvalue of that matrix. Thus, zero is an eigenvalue of $\hat{L}$. The proof of the remaining part is also similar to the corresponding part of Theorem 4.2. We do not go to the details for this too. ∎
Similar to Theorem 5.1, we have the following theorem. The Jordan canonical form is important in the proof.
Theorem 6.3.
Suppose that we have a directional and connected graph $G=(V,E)$ with reasonable desired relative configurations $\left\{\hat{q}_{d_{ij}}:(i,j)\in E\right\}$ on $G$. Let $\hat{\mathbf{q}}_{d}\in\hat{\mathbb{U}}^{n\times 1}$ be a desired formation satisfying (23). Suppose the in-forest dimension of $G$ is one. If there exists a positive diagonal matrix $K\in\mathbb{R}^{n\times n}$ such that $\hat{K}=\hat{Q}^{*}K\hat{Q}$, where $\hat{Q}=diag(\hat{\mathbf{q}}_{d})$, then the observed configuration vector $\hat{\mathbf{z}}=\hat{\mathbf{z}}(t)$ under control law (31) globally asymptotically converges to $\bar{\hat{\mathbf{q}}}_{d}\hat{c}$ with $-^\hat{c}\in{\hat{\mathbb{Q}}}$.
Proof.
From the assumptions we have
$\hat{K}\hat{L}=(\hat{Q}^{*}K\hat{Q})(\hat{Q}^{*}L\hat{Q})=\hat{Q}^{*}KL\hat{Q}.$
Since $K$ is a positive definite diagonal matrix, all eigenvalues of $KL$ have positive real parts except one zero eigenvalue. Denote the Jordan decomposition of $KL$ as $SJS^{-1}$. Without loss of generalization, let $J=diag(0,J_{2})$. Then $J_{2}$ is a Jordan matrix and all diagonal elements have positive real parts. Let $V=cS$. Then there is
$V^{-1}KLV=\begin{bmatrix}0&0\\ 0&J_{2}\end{bmatrix}.$ (38)
Here, $c=\frac{1}{{\bf 1}^{\top}s_{1}}$, $s_{1}$ is the first column of $S$. Therefore, the first column of $V\in\mathbb{R}^{n\times n}$ is $\mathbf{1}_{n}$. Define
${\hat{\mathbf{y}}}=\begin{bmatrix}\hat{y}_{1}\\ {\hat{\mathbf{y}}_{2}}\end{bmatrix}=V^{-1}\hat{Q}\hat{\mathbf{z}}.$
Then we have
$\begin{bmatrix}\dot{\hat{y}}_{1}\\ \dot{{\hat{\mathbf{y}}}}_{2}\end{bmatrix}=-\begin{bmatrix}0&0\\ 0&J_{2}\end{bmatrix}\begin{bmatrix}\hat{y}_{1}\\ {\hat{\mathbf{y}}_{2}}\end{bmatrix}.$ (39)
It follows that $\hat{y}_{1}(t)=\hat{y}_{1}(0)$ and ${\hat{\mathbf{y}}_{2}(t)=\exp(-J_{2}t)\hat{\mathbf{y}}(0)}$ asymptotically converges to a zero vector as $t\rightarrow\infty$. Therefore, $\hat{\mathbf{z}}(t)={\hat{Q}^{*}V\hat{\mathbf{y}}(t)}$ globally asymptotically converges to a formation $\bar{\hat{\mathbf{q}}}_{d}\hat{y}_{1}(0)$. This completes the proof. ∎