### v3

# Dual Quaternion Laplacian, Directed Gain Graph and Formation Control

Abstract
We study the multi-agent formation control problem in a directed graph. We show that a desired relative configuration scheme is reasonable in a directed graph if and only if for any cycle in this directed graph, the product of relative configurations of the forward arcs, and inverses of relative configurations of the backward arcs, is equal to $1$. We then show that for a reasonable desired relative configuration scheme in a directed and connected graph, the dual quaternion Laplacian is similar to the unweighted Laplacian of the directed graph, zero is an eigenvalue of these two Laplacian matrices, and a formation vector is a desired formation satisfying this reasonable desired relative configuration scheme if and only if it is a unit dual quaternion vector and an eigenvector, corresponding to the zero eigenvalue, of the dual quaternion Laplacian matrix. A numerical method is given to determine whether a given desired relative configuration scheme is reasonable or not. A control law is presented when the desired relative configuration scheme is reasonable. Our discussion is extended to general directed gain graphs.
Key words. Dual quaternion Laplacian, directed graph, formation control, relative configuration, directed gain graph.
1 Introduction
The multi-agent formation control problem considers the formation control problem of $n$ rigid bodies in the 3-D space. It is a very important research problem in robotics research. Examples of these $n$ rigid bodies can be autonomous mobile robots, or unmanned aerial vehicles (UAVs), or autonomous underwater vehicles (AUVs), or small satellites. As the unit dual quaternion is a unified mathematical tool to describe the rigid body movement in the 3-D space, dual quaternion is used as a mathematical tool to study the formation control problem [21, 22]. Wang, Yu and Zheng [23] studied the application of dual quaternion matrices in the multi-agent formation control problem. In 2023, Qi, Wang and Luo [17] showed that the relative configuration adjacency matrix and the logarithm adjacency matrix in the multi-agent formation control problem are dual quaternion Hermitian matrices, which have good spectral properties as shown in [16]. They also introduced dual quaternion Laplacian matrices and studied their role in formation control. In multi-agent formation control, a desired relative configuration scheme may be given. People need to know if this scheme is reasonable or not such that a feasible solution of configurations of these multi-agents exists. Recently, by exploring the eigenvalue problem of dual Hermitian matrices, and its link with the unit gain graph theory, Qi and Cui [15] proposed a cross-disciplinary approach to solve this relative configuration problem.
However, as discussed in [22], the multi-agent formation control problem used to occur in a directed graph, i.e., there are some directed control relations among those multi-agents. Hence, in this paper, we consider the multi-agent formation control problem in a directed graph.
In the next section, we review some basic definitions and properties of dual quaternions and dual quaternion matrices.
In Section 3, we study the multi-agent formation control problem in a directed graph. We say a desired relative configuration scheme is reasonable if there is a desired formation vector corresponding to it. We show that a desired relative configuration scheme is reasonable in a directed graph if and only if for any cycle in this directed graph, the product of relative configurations of the forward arcs of the cycle, and inverses of relative configurations of the backward arcs of the cycle, is equal to $1$. We then show that for a reasonable desired relative configuration scheme in a directed and connected graph, the dual quaternion Laplacian is similar to the unweighted Laplacian of the directed graph, zero is an eigenvalue of these two Laplacian matrices, and a formation vector is a desired formation satisfying this reasonable desired relative configuration scheme if and only if it is a a unit dual quaternion vector and an eigenvector, corresponding to the zero eigenvalue, of the dual quaternion Laplacian matrix. Some control laws are also studied for such a directed graph in that section.
Based on the theoretical study in Section 3, in Section 4, we propose a numerical method to determine whether a desired relative configuration scheme on a directed graph is reasonable or not.
Then we present a control law in Section 5, when the desired relative configuration scheme is reasonable.
In Section 6, we extend our study to directed gain graphs. A gain graph is a special directed gain graph, but not vice versa. What we studied in Sections 3 and 4 are dual quaternion directed unit gain graphs. In the literature, there are discussions on signed directed graphs. Directed gain graphs also include complex directed unit gain graphs and quaternion directed unit gain graphs.
2 Dual Quaternions and Dual Quaternion Matrices
2.1 Quaternions and Dual Quaternions
Denote the field of real numbers by $\mathds{R}$, and the space of three-dimensional real vectors by $\mathbb{V}$. Use small letters with overline arrows, such as $\vec{p},\vec{q}$, to denote three-dimensional vectors. For $\vec{p}\in\mathbb{V}$, we denote that $\vec{p}=[p_{1},p_{2},p_{3}]$, but we still regard $\vec{p}$ as a column vector and its transpose $\vec{p}^{\top}$ a row vector. In particular, denote $\vec{0}:=[0,0,0]$.
A quaternion $q=[q_{0},\vec{q}\,]=[q_{0},q_{1},q_{2},q_{3}]$ is a real four-dimensional vector, where $q_{0}\in\mathds{R}$ and $\vec{q}=[q_{1},q_{2},q_{3}]\in\mathbb{V}$. Denote the set of all quaternions by $\mathbb{Q}$. Let $p=[p_{0},\vec{p}\,],q=[q_{0},\vec{q}\,]\in\mathbb{Q}$, where $\vec{p}=[p_{1},p_{2},p_{3}],\vec{q}=[q_{1},q_{2},q_{3}]\in\mathbb{V}$. The sum of quaternions $p$ and $q$ is defined as
$p+q=[p_{0}+q_{0},\vec{p}+\vec{q}\,].$
The product of $p$ and $q$ is defined by
${pq}=[p_{0}q_{0}-\vec{p}\cdot\vec{q},p_{0}\vec{q}+q_{0}\vec{p}+\vec{p}\times\vec{q}\,],$
where $\vec{p}\cdot\vec{q}$ is the dot product, i.e., inner product of $\vec{p}$ and $\vec{q}$, with
$\vec{p}\cdot\vec{q}\equiv\vec{p}^{\top}\vec{q}=p_{1}q_{1}+p_{2}q_{2}+p_{3}q_{3},$
and $\vec{p}\times\vec{q}$ is the cross product of $\vec{p}$ and $\vec{q}$, with
$\vec{p}\times\vec{q}=[p_{2}q_{3}-p_{3}q_{2},-p_{1}q_{3}+p_{3}q_{1},p_{1}q_{2}-p_{2}q_{1}]=-\vec{q}\times\vec{p}.$
Thus, in general, $pq\not=qp$.
The conjugate of a quaternion $q=[q_{0},q_{1},q_{2},q_{3}]\in\mathbb{Q}$ is defined as $q^{*}=[q_{0},-q_{1},-q_{2},-q_{3}]$. For any $p,q\in\mathbb{Q}$, we have
$(pq)^{*}=q^{*}p^{*}.$
Suppose that $q=[q_{0},q_{1},q_{2},q_{3}]\in\mathbb{Q}$. Then its magnitude is defined as
$|q|:=\sqrt{q_{0}^{2}+q_{1}^{2}+q_{2}^{2}+q_{3}^{2}}.$
And $q$ is invertible if and only if $|q|$ is positive. In this case, we have
$q^{-1}={q^{*}\over|q|^{2}}.$ (1)
A quaternion $q\in\mathbb{Q}$ is called a unit quaternion if $|q|=1$. Denote the set of all unit quaternions by $\mathbb{U}$.
If $p,q\in\mathbb{U}$, then $pq\in\mathbb{U}$. For any $q\in\mathbb{U}$,
$\tilde{q}\tilde{q}^{*}=\tilde{q}^{*}\tilde{q}=\tilde{1},$
$q$ is invertible and $q^{-1}=q^{*}$.
A quaternion $q=[0,q_{1},q_{2},q_{3}]\in\mathbb{Q}$ is called a vector quaternion. A quaternion $q\in\mathbb{Q}$ is a vector quaternion if and only if $q=-q^{*}$.
A dual quaternion $\hat{q}=q_{s}+q_{d}\epsilon$ consists of two quaternions $q_{s}$, the standard part of $\hat{q}$, and $q_{d}$, the dual part of $\hat{q}$, with $\epsilon$ as the infinitesimal unit, satisfying $\epsilon\not=0$ but $\epsilon^{2}=0$. If both $q_{s}$ and $q_{d}$ are real numbers, then $\hat{q}$ is a dual number. We use a hat symbol to distinguish a dual quaternion. We denote the set of dual quaternions as $\hat{\mathbb{Q}}$. For $\hat{p}=p_{s}+p_{d}\epsilon\in\hat{\mathbb{Q}}$, if $p_{s}\not=0$, then we say that $\hat{p}$ is appreciable. Let $\hat{p}=p_{s}+p_{d}\epsilon,\hat{q}=q_{s}+q_{d}\epsilon\in\hat{\mathbb{Q}}$. Then the sum of $\hat{p}$ and $\hat{q}$ is
$\hat{p}+\hat{q}=(p_{s}+q_{s})+(p_{d}+q_{d})\epsilon$
and the product of $\hat{p}$ and $\hat{q}$ is
$\hat{p}\hat{q}=p_{s}q_{s}+(p_{s}q_{d}+p_{d}q_{s})\epsilon.$
Again, in general, $\hat{p}\hat{q}\not=\hat{q}\hat{p}$, but for any $\hat{p},\hat{q},\hat{s}\in\hat{\mathbb{Q}}$, we have
$(\hat{p}\hat{q})\hat{s}=\hat{p}(\hat{q}\hat{s}).$
The conjugate of $\hat{p}=p_{s}+p_{d}\epsilon$ is $\hat{p}^{*}=p_{s}^{*}+p_{d}^{*}\epsilon$. The magnitude of $\hat{p}$ is defined as
$|\hat{p}|:=\left\{\begin{aligned} |p_{s}|+{(p_{s}p_{d}^{*}+p_{d}p_{s}^{*})\over 2|p_{s}|}\epsilon,&\ {\rm if}\ p_{s}\not=0,\\ |p_{d}|\epsilon,&\ {\rm otherwise},\end{aligned}\right.$ (2)
which is a dual number.
A dual quaternion $\hat{p}=p_{s}+p_{d}\epsilon$ is called a unit dual quaternion if $p_{s}$ is a unit quaternion and
${p_{s}p_{d}^{*}+p_{d}p_{s}^{*}=0.}$ (3)
The set of unit dual quaternions is denoted as $\hat{\mathbb{U}}$.
A dual quaternion $p=p_{s}+p_{d}\epsilon$ is called a vector dual quaternion if both $p_{s}$ and $p_{d}$ are vector quaternions.
A unit dual quaternion $\hat{q}=q_{s}+q_{d}\epsilon$ can represent the movement of a rigid body as
$\hat{q}=q_{s}+{\epsilon\over 2}q_{s}p^{b},$ (4)
where $q$ is a unit quaternion, representing the rotation of the rigid body, and $p^{b}=[0,\vec{p}\,^{b}]$ is a vector quaternion, with ${\vec{p}\,^{b}}$ representing the translation of the rigid body, and the superscript $b$ relating to the body frame, which is attached to the rigid body [22]. We may use $q$ to represent the configuration of the rigid body. Then $q$ represents the attitude of the rigid body, and $p^{b}$ represents the position of the rigid body.
2.2 Dual Quaternion Vectors and Dual Quaternion Matrices
Denote $\hat{\mathbf{x}}=(\hat{x}_{1},\cdots,\hat{x}_{n})^{\top}\in{\hat{\mathbb{Q}}}^{n}$ for dual quaternion vectors. We say that $\hat{\mathbf{x}}\in\hat{\mathbb{Q}}^{n}$ is appreciable if at least one of its components is appreciable. We denote $\mathbf{x}\in\mathbb{U}^{n}$ if $|x_{i}|=1$ for all $i=1,\dots,n$. We may also write
$\hat{\mathbf{x}}=\mathbf{x}_{s}+\mathbf{x}_{d}\epsilon,$
where $\mathbf{x}_{s},\mathbf{x}_{d}\in{\mathbb{Q}}^{n}$. Define ${\bar{\hat{\mathbf{x}}}}:=(\hat{x}_{1}^{*},\dots,\hat{x}_{n}^{*})^{\top}$ and $\hat{\mathbf{x}}^{*}:=(\hat{x}_{1}^{*},\dots,\hat{x}_{n}^{*})$. The $2$-norm of $\hat{\mathbf{x}}\in\hat{\mathbb{Q}}^{n}$ is defined as
$\|\hat{\mathbf{x}}\|_{2}=\left\{\begin{aligned} \sqrt{\sum_{i=1}^{n}|\hat{x}_{i}|^{2}},&\ {\rm if}\ \hat{\mathbf{x}}\not=\mathbf{0},\\ \|\mathbf{x}_{d}\|_{2}\epsilon,&\ {\rm otherwise.}\end{aligned}\right.$ (5)
If $\hat{\mathbf{x}}^{(1)},\dots,\hat{\mathbf{x}}^{(n)}\in\hat{\mathbb{Q}}^{n}$ and $\left(\hat{\mathbf{x}}^{(i)}\right)^{*}\hat{\mathbf{x}}^{(j)}=\delta_{ij}$ for $i,j=1,\cdots,n$, where $\delta_{ij}$ is the Kronecker symbol, then we say that $\{\hat{\mathbf{x}}^{(1)},\dots,\hat{\mathbf{x}}^{(n)}\}$ is an orthonormal basis of ^$\hat{\mathbb{Q}}^{n}$.
The collection of $m\times n$ dual quaternion matrices is denoted by ^$\hat{\mathbb{Q}}^{m\times n}$.
A dual quaternion matrix $\hat{A}=(\hat{a}_{ij})\in\hat{\mathbb{Q}}^{m\times n}$ can be denoted as
$\hat{A}=A_{s}+A_{d}\epsilon,$ (6)
where $A_{s},A_{d}\in{\mathbb{Q}}^{m\times n}$. We call $A_{s}$ and $A_{d}$ the standard part and the infinitesimal part of $\hat{A}$, respectively. The transpose of $\hat{A}$ is $\hat{A}^{\top}=(\hat{a}_{ji})$. The conjugate transpose of $\hat{A}$ is $\hat{A}^{*}=(\hat{a}_{ji}^{*})$.
Let $\hat{A}\in\hat{\mathbb{Q}}^{m\times n}$ and $\hat{B}\in\hat{\mathbb{Q}}^{n\times r}$. Then we have $(\hat{A}\hat{B})^{*}=\hat{B}^{*}\hat{A}^{*}$.
Given a square dual quaternion matrix $\hat{A}\in\hat{\mathbb{Q}}^{n\times n}$, it is called invertible (nonsingular) if $\hat{A}\hat{B}=\hat{B}\hat{A}=I_{n}$ for some $\hat{B}\in\hat{\mathbb{Q}}^{n\times n}$, where $I_{n}$ is the $n\times n$ identity matrix. Such $\hat{B}$ is unique and denoted by $\hat{A}^{-1}$. Square dual quaternion matrix $\hat{A}$ is called Hermitian if $\hat{A}^{*}=\hat{A}$. Write ${\hat{A}}=A_{s}+A_{d}\epsilon\in\hat{\mathbb{Q}}^{n\times n}$. Then $\hat{A}$ is Hermitian if and only if both $A_{s}$ and $A_{d}$ are quaternion Hermitian matrices. Square dual quaternion matrix $\hat{A}$ is called unitary if $\hat{A}^{*}\hat{A}=I_{n}$. Apparently, $\hat{A}\in\hat{\mathbb{Q}}^{n\times n}$ is unitary if and only if its column vectors form an orthonormal basis of ^$\hat{\mathbb{Q}}^{n}$. The rank of $\hat{A}$ is the number of nonzero singular values of $A$ and the appreciable rank of $\hat{A}$ is the number of appreciable singular values of $A$ [16].
2.3 Eigenvalues of Dual Quaternion Matrices
Suppose that $\hat{A}\in{\hat{\mathbb{Q}}}^{n\times n}$. If there are $\lambda\in\hat{\mathbb{Q}}$, $\hat{\mathbf{x}}\in{\hat{\mathbb{Q}}}^{n}$, where $\hat{\mathbf{x}}$ is appreciable, such that
$\hat{A}\hat{\mathbf{x}}=\hat{\mathbf{x}}\hat{\lambda},$ (7)
then we say that $\hat{\lambda}$ is a right eigenvalue of $\hat{A}$, with $\hat{\mathbf{x}}$ as an associated right eigenvector. If there are $\hat{\lambda}\in\hat{\mathbb{Q}}$, $\hat{\mathbf{x}}\in{\hat{\mathbb{Q}}}^{n}$, where $\hat{\mathbf{x}}$ is appreciable, such that
$\hat{A}\hat{\mathbf{x}}=\hat{\lambda}\hat{\mathbf{x}},$
then we say that $\hat{\lambda}$ is a left eigenvalue of $\hat{A}$, with $\hat{\mathbf{x}}$ as an associated left eigenvector. If $\hat{\lambda}$ is a dual number and a right eigenvalue of $\hat{A}$, then it is also a left eigenvalue of $\hat{A}$, as a dual number is commutative with a dual quaternion vector. In this case, we simply say that it is an eigenvalue of $\hat{A}$, with $\hat{\mathbf{x}}$ as an associated eigenvector.
The theory of eigenvalues of dual quaternion Hermitian matrices, and the theory of singular values of dual quaternion matrices were established in [16]. A recent application of eigenvectors of dual quaternion matrices can be found in [5]. We do not go to details here.
3 Dual Quaternion Laplacian of a Directed Graph
Consider the formation control problem of $n$ rigid bodies, which can be described by a directed (digraph) $G=(V,E)$ with $n$ nodes and $m$ arcs. For two rigid bodies $i$ and $j$ in $V$, if rigid body $i$ can sense rigid body $j$, then arc $(i,j)\in E$, where $i$ and $j$ are the tail and head of the arc $(i,j)$ respectively. It is possible that $(i,j)\in E$ but $(j,i)\not\in E$, i.e., rigid body $i$ can sense rigid body $j$ but rigid body $j$ cannot sense rigid body $i$.
A digraph $G$ is weakly connected or simply called connected if there is an undirected path between any pair of vertices, or equivalently, the underlying undirected graph of $G$ is connected.
Suppose that there is a cycle $C=\{i_{1},i_{2},\dots,i_{k},{i_{k+1}}\equiv i_{1}\}$ in a directed graph $G$. Then $C$ contains $k$ arcs in $E$. Suppose that $1\leq j\leq k$. If arc $\left(i_{j},i_{j+1}\right)\in E$ is a member of $C$, then we say that it is a forward arc in $C$, and denote that $\left(i_{j},i_{j+1}\right)\in C$. If arc $\left(i_{j+1},i_{j}\right)\in E$ is a member of $C$, then we say that it is a backward arc in $C$, and denote $\left(i_{j+1},i_{j}\right)\in C$.
The Laplacian matrix of a directed graph $G$ is defined by
$\hat{L}={D-\hat{A}},$ (8)
where $D$ is a diagonal real matrix whose $i$-th diagonal element is equal to the out-degree $d_{i}$ (the number of arcs coming out from $i$) of the $i$-th vertex, and ${\hat{A}}=(\hat{a}_{ij})$,
$\hat{a}_{ij}=\left\{\begin{array}[]{cl}{\hat{q}_{d_{ij}}},&\text{if }(i,j)\in E,\\ 0,&\text{otherwise},\end{array}\right.$
where $\hat{q}_{d_{ij}}$ is the desired relative configuration on the edge $(i,j)\in E$. Now for $(i,j)\in E$, let $\hat{a}_{ij}\in\hat{\mathbb{U}}$ be the desired relative configuration of the rigid bodies $i$ and $j$. Then $\hat{A}$ is the desired relative configuration adjacency matrix.
Next, we present an example.
Example 3.1 Consider the two figures in Fig. 1. For the the directed tree on the left, we have $V=\{1,2,3\}$, $E=\{(2,1),(3,1)\}$, and
$\hat{A}^{(1)}=\begin{bmatrix}0&0&0\\ \hat{q}_{d_{21}}&0&0\\ \hat{q}_{d_{31}}&0&0\end{bmatrix},\quad\hat{L}^{(1)}=\begin{bmatrix}0&0&0\\ -\hat{q}_{d_{21}}&1&0\\ -\hat{q}_{d_{31}}&0&1\end{bmatrix}.$ (9)
For the directed cycle on the right, we have $V=\{1,2,3\}$, $E=\{(1,3),(2,1),(3,2)\}$, and
$\hat{A}^{(2)}=\begin{bmatrix}0&0&\hat{q}_{d_{13}}\\ \hat{q}_{d_{21}}&0&0\\ 0&\hat{q}_{d_{32}}&0\end{bmatrix},\quad\hat{L}^{(2)}=\begin{bmatrix}1&0&-\hat{q}_{d_{13}}\\ -\hat{q}_{d_{21}}&1&0\\ 0&-\hat{q}_{d_{32}}&1\end{bmatrix}.$ (10)
Figure 1: Two directed grapgs. Left: A directed tree with three vertices. Right: A directed cycle with three vertices.
Note that, for a directed graph, neither $\hat{A}$ nor $\hat{L}$ is a dual quaternion Hermitian matrix. In this case, their spectral analysis is complicated.
The spectral theory of the Laplacian matrix of a directed graph $G$ is important. Let $L$ be the unweighted Laplacian matrix of $G$. The following result from [19] is quite useful in our analysis.
Lemma 3.1 ([19], Corollary 2.5).
The nonsymmetrical Laplacian matrix of a directed graph has a simple zero eigenvalue with an associated eigenvector ${\bf 1}$ and all of the other eigenvalues are in the open right half plane if and only if the directed graph has a directed spanning tree.
Here, $G$ has a directed spanning tree if there exists a node $r$ (a leaf) such that all other nodes can be linked to $r$ via a directed path. In [19], $r$ is a root since the enetries of the adjacency matrix are nonzero if $(j,i)\in E$, which is different from ours.
Note that Lemma 3.1 may not hold for Laplacian matrices of dual quaternion unit gain graphs. For instance, let $\hat{q}_{d_{13}}=[0,1,0,0]$, $\hat{q}_{d_{21}}=[0,0,1,0]$, and $\hat{q}_{d_{32}}=[0,0,0,1]$ in (10), then we have $\mathrm{rank}(\hat{L}^{(2)})=2$. If we change $\hat{q}_{d_{32}}$ to any other unit dual quaternion numbers not equal $[0,0,0,1]$, then we have $\mathrm{rank}(\hat{L}^{(2)})=3$.
Suppose that for each $(i,j)\in E$, we have a desired relative configuration from rigid body $i$ to $j$ as $\hat{q}_{d_{ij}}\in\hat{\mathbb{U}}$. We say that the desired relative configurations $\left\{\hat{q}_{d_{ij}}:(i,j)\in E\right\}$ is reasonable if and only if there is a desired formation $\hat{\mathbf{q}}_{d}\in\hat{\mathbb{U}}^{n\times 1}$, which satisfies
$\hat{q}_{d_{ij}}=\hat{q}_{d_{i}}^{*}\hat{q}_{d_{j}},$ (11)
for all $(i,j)\in E$.
By (11), we have the following result.
Lemma 3.2.
Assume that the desired relative configuration $\left\{\hat{q}_{d_{ij}}:(i,j)\in E\right\}$ is reasonable. Suppose both $(i,j)$ and $(j,i)$ are in $E$. Then we have
$\hat{q}_{d_{ij}}=\hat{q}_{d_{ji}}^{*}.$ (12)
Proof.
This result could be see directed by (11). ∎
The above lemma is only a necessary condition for the desired relative configuration being reasonable. Next, we show an equivalent condition in the following theorem.
Theorem 3.3.
Suppose that we have a directed graph $G=(V,E)$ that has a directed spanning tree. The desired relative configuration $\left\{\hat{q}_{d_{ij}}:(i,j)\in E\right\}$ is reasonable if and only if for any cycle $C=\left\{i_{1},\dots,i_{k},i_{k+1}\equiv i_{1}\right\}$ of $G$, we have
$\prod_{1\leq j\leq k}\hat{p}_{i_{j}i_{j+1}}=1,$ (13)
where $\hat{p}_{i_{j}i_{j+1}}=\hat{q}_{d_{i_{j}i_{j+1}}}$ if $\left(i_{j},i_{j+1}\right)\in C$, and $\hat{p}_{i_{j}i_{j+1}}=\hat{q}_{d_{i_{j}i_{j+1}}}^{*}$ if $\left(i_{j+1},i_{j}\right)\in C$.
Furthermore, let ${\hat{\mathbf{q}}}_{0}$ be a desired formation satisfying (11). Then the set of all the desired formations satisfying (11) is
${[\hat{\mathbf{q}}_{0}]=}\left\{\hat{c}\hat{\mathbf{q}}_{0}:\hat{c}\in\hat{\mathbb{U}}\right\}.$ (14)
Proof.
Suppose that the desired relative configurations $\left\{\hat{q}_{d_{ij}}:(i,j)\in E\right\}$ is reasonable. Then there is a desired formation $\hat{\mathbf{q}}_{d}\in\hat{\mathbb{U}}^{n}$, which satisfies (11) for all $(i,j)\in E$. Then this implies (13).
On the other hand, suppose that the desired relative configurations $\left\{\hat{q}_{d_{ij}}\right\}$ satisfy (13) for all cycles in $G$. Let $m$ be the number of edges of the underlying undirected graph of $G$. Since $G$ is connected, we have $m\geq n-1$. Let $M=m-n+1$. We now show by induction on $M$, that there is a desired formation $\hat{\mathbf{q}}_{d}\in\hat{\mathbb{U}}^{n}$, which satisfies (11) for all $(i,j)\in E$.
(i) We first show this for $M=0,i.e.,m=n-1$. Then $G$ is a tree. Without loss of generality, we pick node $1$ as the root of this tree. Let $\hat{q}_{d_{1}}=\hat{1}$. Then we may determine $\hat{q}_{d_{j}}$ for all $j\not=1$ by (11) from each father node to its children nodes. Then equation (11) is satisfied for all $(i,j)\in E$.
(ii) We now assume that this claim is true for $M=M_{0}$, and prove that it is true for $M=M_{0}+1$. Then $M>m_{0}-n+1$. This implies that there is at least one cycle $\left\{j_{1},\dots,j_{k}\right\}$, with $j_{k+1}=j_{1}$, of $G$, such that the $k$ nodes $j_{1},\dots,j_{k}$ are all distinct and $k\geq 3$. Without loss of generality, we may assume that $(j_{k},j_{1})\in E$. Delete the edge $(j_{k},j_{1})$ from $E$. Let $E_{0}=E\setminus\{(j_{k},j_{1})\}$, and $G_{0}=(V,E_{0})$. Then $G_{0}$ is still directed and connected, and the desired relative configurations $\left\{\hat{q}_{d_{ij}}:(i,j)\in E_{0}\right\}$ satisfy (13) for all cycles in $G_{0}$. By our induction assumption, there is a desired formation $\hat{\mathbf{q}}_{d}\in\hat{\mathbb{U}}^{n\times 1}$ such that (11) holds for all $(i,j)\in E_{0}$. Now, by this and (13) for this cycle $\left\{j_{1},\dots,j_{k}\right\}$, (11) also holds for $(j_{1},j_{k})$. This proves this claim for $M=M_{0}+1$ in this case.
Suppose that ${\hat{\mathbf{q}}}_{0}$ is a desired formation satisfying (11), and $\hat{c}\in\hat{\mathbb{U}}$. Then $\hat{c}{\hat{\mathbf{q}}}_{0}$ also satisfies (11). We may prove the other side of the last claim by induction as above.
The proof is completed. ∎
The set (14) means that while the $n$ rigid bodies converge asymptotically to the desired relative formation, they may still have rotations and translations as a cohesive whole as in [12], or they have a lead which moves following some trajectories and the other rigid bodies follow that lead as in [22].
The equivalent condition in (13) requires enumerating all cycles, which may be expensive for a large scale graph. In the following theorem, we show another equivalent condition based on the zero eigenvalue of the Laplacian matrix and its corresponding eigenvectors.
Theorem 3.4.
Suppose that we have a directed graph $G=(V,E)$ that has a directed spanning tree and there is a reasonable desired relative configuration scheme $\left\{\hat{q}_{d_{ij}}:(i,j)\in E\right\}$ on $G$. Assume that the in-forest dimension of $G$ is one. Let the dual quaternion Laplacian $\hat{L}$ be constructed as above. Then a formation ${\hat{\mathbf{q}}}_{d}$ is a desired formation satisfying (11), if and only if it is in $\hat{\mathbb{U}}^{n}$ and satisfies
$\hat{L}\bar{\hat{\mathbf{q}}}_{d}=\hat{\mathbf{0}},$ (15)
where $\bar{\hat{\mathbf{q}}}_{d}=[{\hat{q}}^{*}_{d_{1}},\cdots,{\hat{q}}^{*}_{d_{n}}]^{\top}$ is the conjugate of $\hat{\mathbf{q}}_{d}$.
Let $\hat{Q}=\mathrm{diag}(\hat{\mathbf{q}}_{d})$. Then $\hat{Q}$ is a unitary matrix, and
$\hat{Q}{\hat{L}}\hat{Q}^{*}=L=D-A,$ (16)
where $D$, $A$, and $L$ are the degree matrix, the unweighted adjacency matrix, and the unweighted Laplacian matrix of $G$.
Furthermore, all eigenvalues of $L$ are right eigenvalues of $\hat{L}$, which are complex numbers. Then zero is an eigenvalue of $\hat{L}$. The conjugates of all the desired formation vectors are eigenvectors of $\hat{L}$, corresponding to the simple zero eigenvalue of $\hat{L}$.
Proof.
Suppose that ${\hat{\mathbf{q}}}_{d}$ is a desired formation. Then it satisfies (11). Then the left of the $i$th equation of (15) satisfies
$d_{i}\hat{q}_{d_{i}}^{*}-\sum_{(i,j)\in E}\hat{q}_{d_{ij}}\hat{q}_{d_{j}}^{*}=d_{i}\hat{q}_{d_{i}}^{*}-\sum_{(i,j)\in E}\hat{q}_{d_{i}}^{*}\hat{q}_{d_{j}}\hat{q}_{d_{j}}^{*}=0.$
Thus, it satisfies (15). From this, we have (16).
Suppose that $\lambda$ is an eigenvalue of $L$ with an eigenvector $\mathbf{x}$. From the definition, we have
$L\mathbf{x}=\lambda\mathbf{x}\Leftrightarrow\hat{Q}\hat{L}\hat{Q}^{*}\mathbf{x}=\lambda\mathbf{x}\Leftrightarrow\hat{L}\hat{Q}^{*}\mathbf{x}=\hat{Q}^{*}\mathbf{x}\lambda\Leftrightarrow\hat{L}\hat{\mathbf{y}}=\hat{\mathbf{y}}\lambda,$
where $\hat{\mathbf{y}}=\hat{Q}^{*}\mathbf{x}$. In other words, all eigenvalues of $L$ are also right eigenvalues of $\hat{L}$. By Lemma 3.1, the Laplacian matrix $L$ has only one zero eigenvalue and its corresponding eigenvector is $\bf 1$. Therefore, zero is also a right eigenvalue of $\hat{L}$ and its corresponding eigenvector is $\hat{Q}^{*}{\bf 1}=\bar{\hat{\mathbf{q}}}_{d}$. By Section 2, a real right eigenvalue of a dual quaternion matrix can be simply called an eigenvalue of that matrix. Thus, zero is an eigenvalue of $\hat{L}$.
On the other hand, suppose that $\hat{\mathbf{q}}\in{\hat{\mathbb{U}}^{n}}$ and satisfies (15). We now need to show that it satisfies (11). Then, by (16), we have ${\hat{L}}=\hat{Q}^{*}L\hat{Q}$ and
$\hat{L}\bar{\hat{\mathbf{q}}}=\hat{\mathbf{0}}\Leftrightarrow\hat{Q}^{*}L\hat{Q}\bar{\hat{\mathbf{q}}}=\hat{\mathbf{0}}\Leftrightarrow L\hat{Q}\bar{\hat{\mathbf{q}}}=\hat{\mathbf{0}}$
By spectral graph theory, the zero eigenvalue of $L$ has an eigenvector $\hat{\bf 1}$. Therefore,
$\hat{Q}\bar{\hat{\mathbf{q}}}={\hat{\bf 1}}\hat{c}$
for any $c\in\hat{\mathbb{Q}}$. Hence,
$\bar{\hat{\mathbf{q}}}=\hat{Q}^{*}{\hat{\bf 1}}\hat{c}=\bar{\hat{\mathbf{q}}}_{d}\hat{c}.$
Therefore, $\hat{\mathbf{q}}=\hat{c}^{*}\hat{\mathbf{q}}_{d}$ and $\hat{c}\in\hat{\mathbb{U}}$. The sufficient condition holds from Theorem 3.3. This also shows that the conjugates of all the desired formation vectors are the eigenvectors of $\hat{L}$, corresponding to the zero eigenvalue of $\hat{L}$. The proof is completed. ∎
This theorem establishes the relation between desired formations and the null space of the dual quaternion Laplacian, and the relation between the dual quaternion Laplacian and the ordinary Laplacian of a directed graph $G$.
4 A Numerical Method to Determine if a Desired Relative Configuration Scheme is Reasonable or Not
Suppose that we have a directed and connected graph $G=(V,E)$, and a desired relative configuration scheme $\left\{\hat{q}_{d_{ij}}:(i,j)\in E\right\}$. A practical question is: Whether such a desired relative configuration scheme is reasonable or not? Based upon the discussion in the last section, we have the following procedure to answer the above question.
1. If there are arcs $(i,j),(j,i)\in E$, check if (12) is satisfied or not. If (12) is not satisfied for such a pair of arcs, then the desired relative configuration scheme is not reasonable.
2. Construct the dual quaternion Laplacian matrix $\hat{L}$ by (32).
3. Solve
$\hat{L}{\hat{\mathbf{x}}}=\hat{\mathbf{0}},\text{ and }\hat{\mathbf{x}}\in\hat{\mathbb{U}}^{n}.$ (17)
If such a solution exists, then by Theorem 3.4, it is a desired formation satisfying (11) and the desired relative configuration scheme is reasonable. Otherwise, it is not.
We may write $\hat{L}=L_{s}+L_{d}\epsilon$ and $\hat{\mathbf{x}}=\mathbf{x}_{s}+\mathbf{x}_{d}\epsilon$. Then (17) is equivalent to two quaternion equations
$L_{s}\mathbf{x}_{s}={\mathbf{0}},$ (18)
where $\mathbf{x}_{s}\in{\mathbb{U}}^{n}$, and
$L_{s}\mathbf{x}_{d}{=-L_{d}\mathbf{x}_{s}},$ (19)
where $x_{sj}x_{dj}^{*}+x_{dj}x_{sj}^{*}=0$ for $j=1,\dots,n$.
However, solving (18) directly may produce a zero solution which does not satisfy $\mathbf{x}_{s}\in{\mathbb{U}}^{n}$, even though the desired relative configuration scheme is reasonable. The situation may be even more complicated when the dimension of the null space of $L_{s}$ is greater than 1. In this case, we must exploit the null space of $L_{s}$ to see whether $\mathbf{x}_{s}\in{\mathbb{U}}^{n}$ exists. In the following, we make the following assumption.
Assumption 4.1.
Given a directed and connected graph $G=(V,E)$, and a desired relative configuration scheme $\left\{\hat{q}_{d_{ij}}:(i,j)\in E\right\}$. Assume the standard part of the Laplacian matrix satisfies $\mathrm{rank}(L_{s})\geq n-1$.
Under Assumption 4.1, either (18) only has zero solution or all solutions of (18) may be denoted by
$[\mathbf{x}_{s}]=\{\mathbf{y}:\mathbf{y}=\mathbf{x}_{s}q,\ q\in\mathbb{Q}\},$ (20)
where $\mathbf{x}_{s}$ is a nonzero solution of (18). Similarly, when Assumption 4.1 holds and $\mathbf{x}_{s}$ satisfies (18), either (19) has no solution at all or all solutions of (19) may be denoted by
$[\mathbf{x}_{d}]=\{\mathbf{y}:\mathbf{y}=\mathbf{x}_{d}-\mathbf{x}_{s}q,\ q\in\mathbb{Q}\},$ (21)
where $\mathbf{x}_{d}$ is a solution of (19).
Different from the real cases, the nonzero solution of (18) is not unique even if the dimension of the null space of $L_{s}$ is equal to 1. In fact, for any $\mathbf{x}_{s}\in{\mathbb{U}}^{n}$ satisfying $L_{s}\mathbf{x}_{s}={\mathbf{0}}$ and any $q\in\mathbb{U}$, $\mathbf{x}_{s}q\in{\mathbb{U}}^{n}$ is also a solution of (18). Thus, we fix $x_{1s}=1$ and solve the reduced system
$L_{2s}\mathbf{x}_{2s}=-L_{1s},$ (22)
where $L_{1s}\in\mathbb{Q}^{n}$ and $L_{2s}\in\mathbb{Q}^{n\times(n-1)}$ are the first and remaining columns of $L_{s}$, respectively, and $\mathbf{x}_{2s}\in\mathbb{Q}^{n-1}$ is the last $n-1$ elements of $\mathbf{x}_{s}$. If (22) does not admit any solutions, then (18) only has zero solution and the configuration scheme is not reasonable. Otherwise, if (22) has a nonzero solution $\mathbf{x}_{s}$, we still have to check whether $\mathbf{x}_{s}\in{\mathbb{U}}^{n}$: If $\mathbf{x}_{s}\in{\mathbb{U}}^{n}$, then we go to next step and compute $\mathbf{x}_{d}$; Otherwise, by the expression in (20) and $x_{1s}=1$, we should choose $q\in\mathbb{U}$, which does not change the magnitude of the entries. Thus, we conclude that $\mathbf{y}\notin{\mathbb{U}}^{n}$ for any $\mathbf{y}\in[\mathbf{x}_{s}]$ and the configuration scheme is not reasonable.
Similarly, the solution of (19) is not unique when (18) has nonzero solutions. Thus, we fix $x_{1d}=0$ and solve the reduced system
$L_{2s}\mathbf{x}_{2d}=-L_{d}\mathbf{x}_{s},$ (23)
where $\mathbf{x}_{2d}\in\mathbb{Q}^{n-1}$ is the last $n-1$ entries of $\mathbf{x}_{d}$. In fact, for any solution $\mathbf{y}_{d}$ of (19), we can set $\mathbf{x}_{d}=\mathbf{y}_{d}-\mathbf{x}_{s}y_{1d}$ and derive $x_{1d}=0$. If (23) is solvable and $x_{sj}x_{dj}^{*}+x_{dj}x_{sj}^{*}=0$ for $j=1,\dots,n$, then the configuration scheme is reasonable. Otherwise, the configuration scheme is not reasonable since for any solution $\mathbf{y}\in[\mathbf{x}_{d}]$, there is
$x_{sj}y_{j}^{*}+y_{j}x_{sj}^{*}=x_{sj}x_{dj}^{*}+x_{dj}x_{sj}^{*}-x_{sj}(q+q^{*})x_{sj}^{*}=x_{sj}x_{dj}^{*}+x_{dj}x_{sj}^{*}-2q_{0},$
where $q_{0}$ is the standard part of $q$. This combining with $x_{1d}=0$ derives $q_{0}=0$. Thus we conclude that for any solution $\mathbf{y}\in[\mathbf{x}_{d}]$, there is $x_{sj}y_{j}^{*}+y_{j}x_{sj}^{*}=x_{sj}x_{dj}^{*}+x_{dj}x_{sj}^{*}$ for $i=1,\dots,n$ and the configuration scheme is not reasonable.
We summarize the whole procedure in the following algorithm.
Algorithm 1 A numerical method to determine if a desired relative configuration scheme is reasonable or not
1:a directed and connected graph $G=(V,E)$, and a desired relative configuration scheme $\left\{\hat{q}_{d_{ij}}:(i,j)\in E\right\}$.
2:If there are arcs $(i,j),(j,i)\in E$, check if (12) is satisfied or not. If (12) is not satisfied for such a pair of arcs, then Flag=0 and stop.
3:Construct the Laplacian matrix $L\in\mathbb{Q}^{n\times n}$ and solve (22).
4:if (22) is solvable and $\mathbf{x}_{s}\in\mathbb{U}^{n}$ then
5:      Solve (23).
6:      if (23) is solvable and $x_{sj}x_{dj}^{*}+x_{dj}x_{sj}^{*}=0$ for $j=1,\dots,n$ then
7:           Flag=1 and stop.
8:      else
9:           Flag=0 and stop.
10:      end if
11:else
12:      Flag=0 and stop.
13:end if
14:Output: Flag.
Let us revisit the examples in Example 3.1. For the example of directed tree, we have
$L_{1s}^{(1)}=\begin{bmatrix}0\\ -q_{d_{21s}}\\ -q_{d_{31s}}\end{bmatrix},\quad L_{2s}^{(1)}=\begin{bmatrix}0&0\\ 1&0\\ 0&1\end{bmatrix},\text{ and }L_{d}^{(1)}=\begin{bmatrix}0&0&0\\ -q_{d_{21d}}&0&0\\ -q_{d_{31d}}&0&0\end{bmatrix}.$
Thus, by solving (22) and (23), we have $\mathbf{x}_{s}^{(1)}=[1,q_{d_{21s}},q_{d_{31s}}]^{\top}$ and $\mathbf{x}_{d}^{(1)}=[0,q_{d_{21d}},q_{d_{31d}}]^{\top}$, respectively. Namely, $\mathbf{x}^{(1)}=[1,\hat{q}_{d_{21}},\hat{q}_{d_{31}}]^{\top}\in\hat{\mathbb{U}}^{3}$ is a solution of (17) and the desired relative configuration scheme is reasonable for any $\hat{q}_{d_{21}},\hat{q}_{d_{31}}\in\hat{\mathbb{U}}$.
For the example of three points directed cycle, we have
$L_{1s}^{(2)}=\begin{bmatrix}1\\ -q_{d_{21s}}\\ 0\end{bmatrix},\quad L_{2s}^{(2)}=\begin{bmatrix}0&-q_{d_{13s}}\\ 1&0\\ -q_{d_{32s}}&1\end{bmatrix},\text{ and }L_{d}^{(2)}=\begin{bmatrix}0&0&-q_{d_{13d}}\\ -q_{d_{21d}}&0&0\\ 0&-q_{d_{32d}}&0\end{bmatrix}.$
Thus, system (22) is solvable only if $q_{d_{13s}}^{*}=q_{d_{32s}}q_{d_{21s}}$. In this case, we have $\mathbf{x}_{s}^{(2)}=[1,q_{d_{21s}},q_{d_{13s}}^{*}]^{\top}$. System (23) is solvable only if $q_{d_{13d}}^{*}=q_{d_{32s}}q_{d_{21d}}+q_{d_{32d}}q_{d_{21s}}$. In this case, we have $\mathbf{x}_{d}^{(2)}=[0,q_{d_{21d}},q_{d_{13d}}^{*}]^{\top}$. Namely, the desired relative configuration scheme is reasonable only if $\hat{q}_{d_{13}}^{*}=\hat{q}_{d_{32}}\hat{q}_{d_{21}}$, which has a solution $\mathbf{x}^{(2)}=[1,\hat{q}_{d_{21}},\hat{q}_{d_{13}}^{*}]^{\top}\in\hat{\mathbb{U}}^{3}$ for (17).
In practice, we may convert (22) and (23) to real or complex equations, whose dimension is expanded to four or two times the original dimension. Then they are surely solvable. For instance, any quaternion linear equation $A\mathbf{x}=\mathbf{b}$ with $A\in\mathbb{Q}^{m\times n}$, $\mathbf{b}\in\mathbb{Q}^{m}$ can be equivalently rewritten to a real matrix equation
$\mathcal{R}(A)\mathcal{R}_{v}(\mathbf{x})=\mathcal{R}_{v}(\mathbf{b}),$ (24)
where $\mathcal{R}(\cdot)$ is a linear homeomorphic mapping from quaternion matrices (or vectors) to their real counterpart and $\mathcal{R}_{v}(\cdot)$ denotes the first column of $\mathcal{R}(\cdot)$. Specifically, let $A=A_{0}+A_{1}i+A_{2}j+A_{3}k$, then
$\mathcal{R}(A)=\begin{bmatrix}A_{0}&-A_{1}&-A_{2}&-A_{3}\\ A_{1}&A_{0}&-A_{3}&A_{2}\\ A_{2}&A_{3}&A_{0}&-A_{1}\\ A_{3}&-A_{2}&A_{1}&A_{0}\end{bmatrix}\text{ and }\mathcal{R}_{v}(A)=\begin{bmatrix}A_{0}\\ A_{1}\\ A_{2}\\ A_{3}\end{bmatrix}$
Several structure-preserving algorithms such as [4, 10, 11] can be used here to solve (24) and (23) efficiently.
5 Control Law
Now suppose that desired relative configuration scheme is reasonable. In this section, we propose the following control law
$\dot{\hat{\mathbf{z}}}=-\hat{K}\hat{L}\hat{\mathbf{z}},$ (25)
or equivalently,
$\dot{\hat{z}}_{i}=\hat{k}_{i}\left(\sum_{j\in\mathcal{N}_{i}}{\hat{q}_{d_{ij}}}\hat{z}_{j}-d_{i}\hat{z}_{i}\right)$ (26)
for $i=1,\dots,n$. Then the following theorem show the asymptotically convergence of our proposed control law. The Jordan canonical form is important in the proof.
Theorem 5.1.
Suppose that we have a directed graph $G=(V,E)$ that has a directed spanning tree and there is a reasonable desired relative configuration $\left\{\hat{q}_{d_{ij}}:(i,j)\in E\right\}$ on $G$. Let $\hat{\mathbf{q}}_{d}\in\hat{\mathbb{U}}^{n\times 1}$ be a desired formation satisfying (11). Suppose the in-forest dimension of $G$ is one. If there exists a positive diagonal matrix $K\in\mathbb{R}^{n\times n}$ such that $\hat{K}=\hat{Q}^{*}K\hat{Q}$, where $\hat{Q}=diag(\hat{\mathbf{q}}_{d})$, then the observed configuration vector $\hat{\mathbf{z}}=\hat{\mathbf{z}}(t)$ under control law (25) globally asymptotically converges to $\bar{\hat{\mathbf{q}}}_{d}\hat{c}$ with $\hat{c}\in{\hat{\mathbb{Q}}}$.
Proof.
From the assumptions we have
$\hat{K}\hat{L}=(\hat{Q}^{*}K\hat{Q})(\hat{Q}^{*}L\hat{Q})=\hat{Q}^{*}KL\hat{Q}.$
Since $K$ is a positive definite diagonal matrix, all eigenvalues of $KL$ have positive real parts except one zero eigenvalue. Denote the Jordan decomposition of $KL$ as $SJS^{-1}$. Without loss of generalization, let $J=diag(0,J_{2})$. Then $J_{2}$ is a Jordan matrix and all diagonal elements have positive real parts. Let $V=cS$. Then there is
$V^{-1}KLV=\begin{bmatrix}0&0\\ 0&J_{2}\end{bmatrix}.$ (27)
Here, $c=\frac{1}{{\bf 1}^{\top}s_{1}}$, $s_{1}$ is the first column of $S$. Therefore, it follows from the fact that the eigenvector corresponding to the zero eigenvalue of $L$ is $\mathbf{1}_{n}$ that the first column of $V\in\mathbb{R}^{n\times n}$ is $\mathbf{1}_{n}$. Define
${\hat{\mathbf{y}}}=\begin{bmatrix}\hat{y}_{1}\\ {\hat{\mathbf{y}}_{2}}\end{bmatrix}=V^{-1}\hat{Q}\hat{\mathbf{z}}.$
Then we have
$\begin{bmatrix}\dot{\hat{y}}_{1}\\ \dot{{\hat{\mathbf{y}}}}_{2}\end{bmatrix}=-\begin{bmatrix}0&0\\ 0&J_{2}\end{bmatrix}\begin{bmatrix}\hat{y}_{1}\\ {\hat{\mathbf{y}}_{2}}\end{bmatrix}.$ (28)
It follows that $\hat{y}_{1}(t)=\hat{y}_{1}(0)$ and ${\hat{\mathbf{y}}_{2}(t)=\exp(-J_{2}t)\hat{\mathbf{y}}(0)}$ asymptotically converges to a zero vector as $t\rightarrow\infty$. Therefore, $\hat{\mathbf{z}}(t)={\hat{Q}^{*}V\hat{\mathbf{y}}(t)}$ globally asymptotically converges to a formation $\bar{\hat{\mathbf{q}}}_{d}\hat{y}_{1}(0)$. This completes the proof.
Now we draw a remark on the relationship between the control law in (25) and several state-of-the-art methods.
Remark 5.1. The graph Laplacian has a long history in the control law design, especially in the consensus problem. In 2004, Olfati-Saber and Murray [14] considered the linear system
$\dot{\mathbf{x}}(t)=-L\mathbf{x}(t),$ (29)
where $L$ is the graph Laplacian induced by the weighted digraph $G=(V,E,A)$ and $A$ is the weighted adjacency matrix. They showed that if $G$ is a connected undirected graph or a strongly connected digraph, then the algorithm in (29) asymptotically solves a consensus problem for all initial states. Furthermore, Olfati-Saber and Murray [14] established a direct connection between the algebraic connectivity of the network and the performance of the control. In 2005, Ren and Beard [18] consided consensus seeking in multi-agent systems under dynamically changing interaction topologies. In 2007, Olfati-Saber, Fax and Murray [13] provided a theoretical framework for analysis of consensus algorithms for multi-agent networked systems.
In 2020, Savino et al. [20] presented a solution based on dual quaternion algebra to the general problem of pose consensus for systems composed of multiple rigid-bodies, i.e.,
$\dot{\hat{\mathbf{y}}}=-L\hat{\mathbf{y}},$ (30)
where $\hat{\mathbf{y}}\in\hat{\mathbb{Q}}^{n}$ is a pure dual quaternion vector. They further showed the proposed pose consensus protocol has guaranteed convergence when the interaction among the agents is represented by directed graphs with directed spanning trees.
Our control law (25) can be regarded as a generalization of (29) in [14] and (30) in [20] from the consensus problem to a general formation control problem. In fact, for a consensus problem, we have $\hat{q}_{d_{1}}=\dots=\hat{q}_{d_{n}}$ and $\hat{q}_{d_{ij}}=1$ for all $(i,j)\in E$. Thus, there is $\hat{L}=L$. In this case, by setting $\hat{K}=I$, (25) reduces to $\dot{\hat{\mathbf{z}}}=-L\hat{\mathbf{z}}.$ This is exactly the control law (30) in [20].
6 Directed Gain Graphs
The discussion in Sections 3 and 4 based on unit dual quaternions may be extended to more general mathematical groups.
Suppose that we have a directed graph $G=(V,E)$, a gain group $\Omega$ and a gain function $\varphi:E\to\Omega$. Then we say that $\Phi=(G,\Omega,\varphi)$ is a directed gain graph. Here, $\Omega$ should be consistent with the real numbers. If $\Omega$ is the group of unit numbers of a number system, then $\Phi$ is called a directed unit gain graph. Let arc $(i,j)\in E$, and $\varphi(i,j)=q_{ij}\in\Omega$. We call $q_{ij}$ the gain on the arc $(i,j)$. In the last two sections, we have $\Omega=\hat{\mathbb{U}}$. We now allow $\Omega$ to be a general mathematical group.
Such a directed gain graph is more general than gain graphs [3, 6, 7, 8, 9, 25]. Actually, a gain graph is a directed gain graph with two additional requirements: 1. For each $(i,j)\in E$, we have $(j,i)\in E$; 2. For each $(i,j)\in E$, we have $q_{ji}=q_{ij}^{-1}$. The discussion in the last two sections shows that general directed gain graphs arise from applications. It also shows that there are some challenging issues there. When $\Omega=\hat{\mathbb{U}}$, the adjacency matrix $A$ and the Laplacian matrix are not Hermitian matrices, while there is no theory on eigenvalues of non-Hermitian dual quaternion matrices.
A walk $W$ in a directed gain graph $\Phi$ is defined in this way. It consists of vertices $i_{1},i_{2},\dots,i_{k}$. For $j=1,\dots,k-1$, either $(i_{j},i_{j+1})\in W$ or $(i_{j+1},i_{j})\in W$. If $(i_{j},i_{j+1})\in W$, we say that the arc $(i_{j},i_{j+1})$ is a forward arc of $W$ and let the shadow element $p_{i_{j}i_{j+1}}=q_{i_{j}i_{j+1}}$. If $(i_{j+1},i_{j})\in W$, we say that the arc $(i_{j+1},i_{j})$ is a backward arc of $W$ and let the shadow element $p_{i_{j}i_{j+1}}=q_{i_{j+1}i_{j}}^{-1}$. The gain of the walk $W$ is
$\varphi(W)=\prod_{j=1}^{k-1}p_{i_{j}i_{j+1}}.$ (31)
The walk $W$ is neutral if $\varphi(W)=1_{\Omega}$, where $1_{\Omega}$ is the identity of $\Omega$. An arc set $S\subseteq E$ is balanced if every cycle $C\subseteq S$ is neutral.
Let $A=(a_{ij})$,
$a_{ij}=\left\{\begin{array}[]{cl}{q_{ij}},&\text{if }(i,j)\in E,\\ 0,&\text{otherwise},\end{array}\right.$
The matrix $A=(a_{ij})$ is called the adjacency matrix of $\Phi$. The Laplacian matrix of a directed gain graph $\Phi$ is defined by
$\hat{L}={D-\hat{A}},$ (32)
where $D$ is a diagonal real matrix whose $i$-th diagonal element is equal to the degree $d_{i}$ of the $i$-th vertex.
A potential function is a function $\theta:V\to\Omega$ such that for every arc $(i,j)\in E$,
$\varphi(i,j)=\theta(i)^{-1}\theta(j).$ (33)
A switching function is a function $\zeta:V\to\Omega$ that switches the directed gain graph $\Phi=(G,\Omega,\varphi)$ to $\Phi^{\zeta}=(G,\Omega,\varphi^{\zeta})$, where
$\varphi^{\zeta}(i,j)=(\zeta(i))^{-1}\varphi(i,j)\zeta(j).$ (34)
Thus, in this case, $\Phi$ and $\Phi^{\zeta}$ are switching equivalent, denoted by $\Phi\sim\Phi^{\zeta}$. Denote $1_{\Omega}$ as a gain mapping such that for any $(i,j)\in E$, we have $1_{\Omega}(i,j)\equiv 1_{\Omega}$. Then we have the following theorem.
Theorem 6.1.
Let $\Phi=(G,\Omega,\varphi)$ be a directed gain graph, which has a directed spanning tree. Then the following are equivalent.
(i) $\Phi$ is balanced.
(ii) $\Phi\sim(G,\Omega,1_{\Omega})$.
(iii) $\varphi$ has a potential function.
Proof.
The proof of the equivalence between (i) and (iii) is similar to the proof of Theorem 3.3.
Suppose that (ii) holds. Let $\varphi^{\zeta}=1_{\Omega}$ and $\theta=\left(\zeta\right)^{-1}$ in (34). Then we have (33), i.e., (iii) holds.
On the other hand, suppose that (iii) holds. Then we have (33) for a potential function $\theta$. Letting $\zeta=\theta^{-1}$ and substituting (33) to (34), we have $\varphi^{\zeta}=1_{\Omega}$, i.e., (ii) holds.
This completes the proof. ∎
We may also extend the other results in Section 3 to general directed gain graphs.
If $\Omega$ is a set of unit numbers, then $\Phi=(G,\Omega,\varphi)$ is called a directed unit gain graph. The first example is $\Omega=\{-1,1\}$, the set of unit real numbers. In the literature, there are called signed directed graphs [24]. Then we may have $\Omega={\mathbb{T}}$, the set of unit complex numbers, or $\Omega={\mathbb{U}}$, the set of unit quaternions, or $\Omega=\hat{\mathbb{U}}$, the set of unit dual quaternions. We may call $\Phi=(G,{\mathbb{T}},\varphi)$ a complex directed unit gain graph, $\Phi=(G,{\mathbb{U}},\varphi)$ a quaternion directed unit gain graph, and $\Phi=(G,\hat{\mathbb{U}},\varphi)$ a dual quaternion directed unit gain graph. What we discussed in the last two sections are dual quaternion directed unit gain graphs.