### v5

# Dual Quaternion Weighted Directed Graph and Formation Control

Abstract
We first study the multi-agent formation control problem in a directed graph. The relative configurations are expressed by unit dual quaternions (UDQs). We call such a weighted directed graph a unit dual quaternion weighted directed graph (UDQWDG). We show that a desired relative configuration scheme is reasonable in a UDQWDG if and only if for any cycle in this directed graph, the product of relative configurations of the forward arcs, and inverses of relative configurations of the backward arcs, is equal to $1$. We then show that a desired relative configuration scheme in a directed connected graph is reasonable if and only if the dual quaternion Laplacian is similar to the unweighted Laplacian of the directed graph. Then for a reasonable desired relative configuration scheme, we build the relationship between the desired formation and the eigenvector corresponding to the zero eigenvalue. A numerical method and a control law are presented. We then study dual quaternion weighted directed graphs (DQWDG). Ordinary graphs, gain graphs, signed directed graphs, complex weighted directed graphs and UDQWDGs are special cases of DQWDGs. A general theory of DQWDG is presented.
Key words. Unit dual quaternion, weighted directed graph, formation control, relative configuration, Non-positive cycle.
1 Introduction
The multi-agent formation control problem considers the formation control problem of $n$ rigid bodies in the 3-D space. It is a very important research problem in robotics research. Examples of these $n$ rigid bodies can be autonomous mobile robots, or unmanned aerial vehicles (UAVs), or autonomous underwater vehicles (AUVs), or small satellites. As the unit dual quaternion is a unified mathematical tool to describe the rigid body movement in the 3-D space, dual quaternion is used as a mathematical tool to study the formation control problem [26, 27]. In multi-agent formation control, a desired relative configuration scheme may be given. People need to know if this scheme is reasonable or not such that a feasible solution of configurations of these multi-agents exists. Recently, by exploring the eigenvalue problem of dual Hermitian matrices, and its link with the unit gain graph theory, Qi and Cui [18] proposed a cross-disciplinary approach to solve this relative configuration problem. Cui, Lu, Qi and Wang [4] investigated dual quaternion unit gain graphs. On the other hand, Chen, Hu, Wang and Guo [3] investigated the application of dual quaternion matrices and dual quaternion unit gain graphs, to the precise formation flying problem of satellite clusters.
However, as discussed in [27], the multi-agent formation control problem used to occur in a directed graph, i.e., there are some directed control relations among those multi-agents. Hence, in Sections 2-4 of this paper, we consider the multi-agent formation control problem in a directed graph.
In the next section, we study the multi-agent formation control problem in a directed graph. Such a directed graph is weighted by unit dual quaternions. Thus, we call such a directed graph a unit dual quaternion weighted directed graph (UDQWDG). We say a desired relative configuration scheme is reasonable if there is a desired formation vector corresponding to it. We show that a desired relative configuration scheme is reasonable in a UDQWDG if and only if for any cycle in this UDQWDG, the product of relative configurations of the forward arcs of the cycle, and inverses of relative configurations of the backward arcs of the cycle, is equal to $1$. We then show that a desired relative configuration scheme in a connected UDQWDG is reasonable if and only if the dual quaternion Laplacian is similar to the unweighted Laplacian of the directed graph. Suppose that the desired relative configuration scheme is reasonable, then zero is an eigenvalue of these two Laplacian matrices, and if in addition, the underlying graph has a directed spanning tree, then a formation vector is a desired formation satisfying this reasonable desired relative configuration scheme if and only if it is a a unit dual quaternion vector and an eigenvector, corresponding to the zero eigenvalue, of the dual quaternion Laplacian matrix.
In Section 3, we propose a numerical method to determine whether a desired relative configuration scheme on a directed graph is reasonable or not.
Then we present a control law in Section 4, when the desired relative configuration scheme is reasonable.
The study above motivates us to consider a general frame of dual quaternion weighted directed graphs (DQWDG) in Section 5. Ordinary graphs, gain graphs, signed directed graphs, complex weighted directed graphs and UDQWDG are special cases of DQWDGs. A general theory for DQWDGs is presented in Section 5.
2 Unit Dual Quaternion Weighted Directed Graph
Consider the formation control problem of $n$ rigid bodies, which can be described by a directed (digraph) $G=(V,E)$ with $n$ nodes and $m$ arcs. For two rigid bodies $i$ and $j$ in $V$, if rigid body $i$ can sense rigid body $j$, then arc $(i,j)\in E$, where $i$ and $j$ are the tail and head of the arc $(i,j)$ respectively. It is possible that $(i,j)\in E$ but $(j,i)\not\in E$, i.e., rigid body $i$ can sense rigid body $j$ but rigid body $j$ cannot sense rigid body $i$.
A digraph $G$ is weakly connected or simply called connected if there is an undirected path between any pair of vertices, or equivalently, the underlying undirected graph of $G$ is connected.
Suppose that there is a cycle $C=\{i_{1},i_{2},\dots,i_{k},{i_{k+1}}\equiv i_{1}\}$ in a directed graph $G$. Then $C$ contains $k$ arcs in $E$. Suppose that $1\leq j\leq k$. If arc $\left(i_{j},i_{j+1}\right)\in E$ is a member of $C$, then we say that it is a forward arc in $C$, and denote that $\left(i_{j},i_{j+1}\right)\in C$. If arc $\left(i_{j+1},i_{j}\right)\in E$ is a member of $C$, then we say that it is a backward arc in $C$, and denote $\left(i_{j+1},i_{j}\right)\in C$.
Denote the conjugate of a quaternion $q$ by $q^{*}$. Let $\mathbb{Q}$ be the ring of quaternions. A dual quaternion $\hat{q}=q_{s}+q_{d}\epsilon$ consists of two quaternions $q_{s}$, the standard part of $\hat{q}$, and $q_{d}$, the dual part of $\hat{q}$, with $\epsilon$ as the infinitesimal unit, satisfying $\epsilon\not=0$ but $\epsilon^{2}=0$. If both $q_{s}$ and $q_{d}$ are real numbers, then the dual quaternion $\hat{q}$ is called a dual number. The magnitude of a dual quaternion $\hat{q}=q_{s}+q_{d}\epsilon$ is defined as
$|\hat{q}|:=\left\{\begin{aligned} |q_{s}|+{(q_{s}q_{d}^{*}+q_{d}q_{s}^{*})\over 2|q_{s}|}\epsilon,&\ {\rm if}\ q_{s}\not=0,\\ |q_{d}|\epsilon,&\ {\rm otherwise},\end{aligned}\right.$ (1)
which is a dual number. A dual quaternion $\hat{q}=q_{s}+q_{d}\epsilon$ is called a unit dual quaternion (UDQ) if $|\hat{q}|=1$, i.e., $q_{s}$ is a unit quaternion and $q_{s}q_{d}^{*}+q_{d}q_{s}^{*}=0$. Let $\hat{\mathbb{Q}}$ be the ring of quaternions. The set of unit dual quaternions is denoted as $\hat{\mathbb{U}}$. A UDQ $\hat{q}\in\hat{\mathbb{U}}$ can represent the movement of a rigid body as
$\hat{q}=q_{s}+{\epsilon\over 2}q_{s}p^{b},$ (2)
where $q_{s}$ is a unit quaternion, representing the rotation of the rigid body, and $p^{b}=[0,\vec{p}\,^{b}]$ and ${\vec{p}\,^{b}}$ representing the translation of the rigid body, and the superscript $b$ relating to the body frame, which is attached to the rigid body [27]. We may use $\hat{q}$ to represent the configuration of the rigid body. Then $q_{s}$ represents the attitude of the rigid body, and $p^{b}$ represents the position of the rigid body. For more knowledge on quaternions and dual quaternions, see [20, 27].
In the formation control problem of $n$ rigid bodies, the related configuration of rigid body are also represented by UDQs.
Suppose that for each $(i,j)\in E$, we have a desired relative configuration from rigid body $i$ to $j$ as a UDQ $\hat{q}_{d_{ij}}\in\hat{\mathbb{U}}$. We say that the desired relative configuration scheme $\left\{\hat{q}_{d_{ij}}:(i,j)\in E\right\}$ is reasonable if and only if there is a desired formation $\hat{\mathbf{q}}_{d}\in{\hat{\mathbb{U}}^{n}}$, which satisfies
$\hat{q}_{d_{ij}}=\hat{q}_{d_{i}}^{*}\hat{q}_{d_{j}},$ (3)
for all $(i,j)\in E$.
We may consider $\hat{q}_{d_{ij}}$ as a weight of the arc $(i,j)$. Then we have a UDQ weighted directed graph (WDG).
The Laplacian matrix of a UDQWDG $G$ is defined by
$\hat{L}={D-\hat{A}},$ (4)
where $D$ is a diagonal real matrix whose $i$-th diagonal element is equal to the out-degree $d_{i}$ (the number of arcs going out from $i$) of the $i$-th vertex: and ${\hat{A}}=(\hat{a}_{ij})$,
$\hat{a}_{ij}=\left\{\begin{array}[]{cl}{\hat{q}_{d_{ij}}},&\text{if }(i,j)\in E,\\ 0,&\text{otherwise},\end{array}\right.$
where $\hat{q}_{d_{ij}}$ is the desired relative configuration on the edge $(i,j)\in E$. Now for $(i,j)\in E$, let $\hat{a}_{ij}\in\hat{\mathbb{U}}$ be the desired relative configuration of the rigid bodies $i$ and $j$. Then $\hat{A}$ is the desired relative configuration adjacency matrix.
Next, we present an example.
Example 2.1.
Consider the two figures in Fig. 1. For the the directed tree on the left, we have $V=\{1,2,3\}$, $E=\{(2,1),(3,1)\}$, and
$\hat{A}^{(1)}=\begin{bmatrix}0&0&0\\ \hat{q}_{d_{21}}&0&0\\ \hat{q}_{d_{31}}&0&0\end{bmatrix},\quad\hat{L}^{(1)}=\begin{bmatrix}0&0&0\\ -\hat{q}_{d_{21}}&1&0\\ -\hat{q}_{d_{31}}&0&1\end{bmatrix}.$ (5)
For the directed cycle on the right, we have $V=\{1,2,3\}$, $E=\{(1,3),(2,1),(3,2)\}$, and
$\hat{A}^{(2)}=\begin{bmatrix}0&0&\hat{q}_{d_{13}}\\ \hat{q}_{d_{21}}&0&0\\ 0&\hat{q}_{d_{32}}&0\end{bmatrix},\quad\hat{L}^{(2)}=\begin{bmatrix}1&0&-\hat{q}_{d_{13}}\\ -\hat{q}_{d_{21}}&1&0\\ 0&-\hat{q}_{d_{32}}&1\end{bmatrix}.$ (6)
Figure 1: Two directed graphs. Left: A directed tree with three vertices. Right: A directed cycle with three vertices.
Note that, for a directed graph, neither $\hat{A}$ nor $\hat{L}$ is a dual quaternion Hermitian matrix. In this case, their spectral analysis is complicated.
The spectral theory of the Laplacian matrix of a directed graph $G$ is important. Let $L$ be the unweighted Laplacian matrix of $G$. The following result from [23] is quite useful in our analysis.
Lemma 2.2 ([23], Corollary 2.5).
The nonsymmetrical Laplacian matrix of a directed graph has a simple zero eigenvalue with an associated eigenvector ${\bf 1}$ and all of the other eigenvalues are in the open right half plane if and only if the directed graph has a directed spanning tree.
Here, $G$ has a directed spanning tree if there exists a node $r$ (a leaf) such that all other nodes can be linked to $r$ via a directed path. In [23], $r$ is a root since the entries of the adjacency matrix are nonzero if $(j,i)\in E$, which is different from ours.
Note that Lemma 2.2 may not hold for Laplacian matrices of dual quaternion unit gain graphs. For instance, let $\hat{q}_{d_{13}}=[0,1,0,0]$, $\hat{q}_{d_{21}}=[0,0,1,0]$, and $\hat{q}_{d_{32}}=[0,0,0,1]$ in (6), then we have $\mathrm{rank}(\hat{L}^{(2)})=2$. If we change $\hat{q}_{d_{32}}$ to any other unit dual quaternion numbers not equal $[0,0,0,1]$, then we have $\mathrm{rank}(\hat{L}^{(2)})=3$.
By (3), we have the following result.
Lemma 2.3.
Assume that the desired relative configuration $\left\{\hat{q}_{d_{ij}}:(i,j)\in E\right\}$ is reasonable. Suppose both $(i,j)$ and $(j,i)$ are in $E$. Then we have
$\hat{q}_{d_{ij}}=\hat{q}_{d_{ji}}^{*}.$ (7)
Proof.
This result could be seen directly by (3). ∎
The above lemma is only a necessary condition for the desired relative configuration scheme being reasonable. Next, we show an equivalent condition in the following theorem.
Theorem 2.4.
Suppose that we have a UDQWDG $G=(V,E)$ that has a directed spanning tree. The desired relative configuration $\left\{\hat{q}_{d_{ij}}:(i,j)\in E\right\}$ is reasonable if and only if for any cycle $C=\left\{i_{1},\dots,i_{k},i_{k+1}\equiv i_{1}\right\}$ of $G$, we have
$\prod_{1\leq j\leq k}\hat{p}_{i_{j}i_{j+1}}=1,$ (8)
where $\hat{p}_{i_{j}i_{j+1}}=\hat{q}_{d_{i_{j}i_{j+1}}}$ if $\left(i_{j},i_{j+1}\right)\in E$, and $\hat{p}_{i_{j}i_{j+1}}=\hat{q}_{d_{i_{j}i_{j+1}}}^{*}$ if $\left(i_{j+1},i_{j}\right)\in E$.
Furthermore, let ${\hat{\mathbf{q}}}_{0}$ be a desired formation satisfying (3). Then the set of all the desired formations satisfying (3) is
${[\hat{\mathbf{q}}_{0}]=}\left\{\hat{c}\hat{\mathbf{q}}_{0}:\hat{c}\in\hat{\mathbb{U}}\right\}.$ (9)
Proof.
Suppose that the desired relative configurations $\left\{\hat{q}_{d_{ij}}:(i,j)\in E\right\}$ is reasonable. Then there is a desired formation $\hat{\mathbf{q}}_{d}\in\hat{\mathbb{U}}^{n}$, which satisfies (3) for all $(i,j)\in E$. Then this implies (8).
On the other hand, suppose that the desired relative configurations $\left\{\hat{q}_{d_{ij}}\right\}$ satisfy (8) for all cycles in $G$. Let $m$ be the number of edges of the underlying undirected graph of $G$. Since $G$ is connected, we have $m\geq n-1$. Let $M=m-n+1$. We now show by induction on $M$, that there is a desired formation $\hat{\mathbf{q}}_{d}\in\hat{\mathbb{U}}^{n}$, which satisfies (3) for all $(i,j)\in E$.
(i) We first show this for $M=0,i.e.,m=n-1$. Then $G$ is a tree. Without loss of generality, we pick node $1$ as the root of this tree. Let $\hat{q}_{d_{1}}=\hat{1}$. Then we may determine $\hat{q}_{d_{j}}$ for all $j\not=1$ by (3) from each father node to its children nodes. Then equation (3) is satisfied for all $(i,j)\in E$.
(ii) We now assume that this claim is true for $M=M_{0}$, and prove that it is true for $M=M_{0}+1$. Then $M>m_{0}-n+1$. This implies that there is at least one cycle $\left\{j_{1},\dots,j_{k}\right\}$, with $j_{k+1}=j_{1}$, of $G$, such that the $k$ nodes $j_{1},\dots,j_{k}$ are all distinct and $k\geq 3$. Without loss of generality, we may assume that $(j_{k},j_{1})\in E$. Delete the edge $(j_{k},j_{1})$ from $E$. Let $E_{0}=E\setminus\{(j_{k},j_{1})\}$, and $G_{0}=(V,E_{0})$. Then $G_{0}$ is still directed and connected, and the desired relative configurations $\left\{\hat{q}_{d_{ij}}:(i,j)\in E_{0}\right\}$ satisfy (8) for all cycles in $G_{0}$. By our induction assumption, there is a desired formation $\hat{\mathbf{q}}_{d}\in\hat{\mathbb{U}}^{n\times 1}$ such that (3) holds for all $(i,j)\in E_{0}$. Now, by this and (8) for this cycle $\left\{j_{1},\dots,j_{k}\right\}$, (3) also holds for $(j_{1},j_{k})$. This proves this claim for $M=M_{0}+1$ in this case.
Suppose that ${\hat{\mathbf{q}}}_{0}$ is a desired formation satisfying (3), and $\hat{c}\in\hat{\mathbb{U}}$. Then $\hat{c}{\hat{\mathbf{q}}}_{0}$ also satisfies (3). We may prove the other side of the last claim by induction as above.
The proof is completed. ∎
The set (9) means that while the $n$ rigid bodies converge asymptotically to the desired relative formation, they may still have rotations and translations as a cohesive whole as in [14], or they have a lead which moves following some trajectories and the other rigid bodies follow that lead as in [27].
The equivalent condition in (8) requires enumerating all cycles, which may be expensive for a large scale graph. In the following theorem, we show another equivalent condition based on the zero eigenvalue of the Laplacian matrix and its corresponding eigenvectors. For the knowledge of eigenvalues of a dual quaternion matrix, see [19].
Theorem 2.5.
Suppose that we have a UDQWDG $\Phi=(G,\Omega,\varphi)$, where $G=(V,E)$, and there is a desired relative configuration scheme $\left\{\hat{q}_{d_{ij}}:(i,j)\in E\right\}$ on $G$. Let the dual quaternion Laplacian $\hat{L}$ be constructed as above. Then there is a desired formation ${\hat{\mathbf{q}}}_{d}$ satisfying (3), if and only if $\hat{\mathbf{q}}_{d}\in\hat{\mathbb{U}}^{n}$ and satisfies
$\hat{Q}{\hat{L}}\hat{Q}^{*}=L=D-A,$ (10)
where $\hat{Q}=\mathrm{diag}(\hat{\mathbf{q}}_{d})$, $D$, $A$, and $L$ are the out-degree matrix, the unweighted adjacency matrix, and the unweighted Laplacian matrix of $G$, respectively.
Furthermore, assume that (10) holds. Then all eigenvalues of $L$ are right eigenvalues of $\hat{L}$, which are complex numbers and
$\hat{L}\bar{\hat{\mathbf{q}}}_{d}=\hat{\mathbf{0}},$ (11)
where $\bar{\hat{\mathbf{q}}}_{d}=[{\hat{q}}^{*}_{d_{1}},\cdots,{\hat{q}}^{*}_{d_{n}}]^{\top}$ is the conjugate of $\hat{\mathbf{q}}_{d}$. In other words, zero is an eigenvalue of $\hat{L}$ and the conjugates of all the desired formation vectors are eigenvectors of $\hat{L}$, corresponding to the zero eigenvalue of $\hat{L}$.
In addition, if $G$ has a directed spanning tree, then a dual quaternion vector ${\hat{\mathbf{q}}}_{d}$ is a desired formation satisfying (3), if and only if it is in $\hat{\mathbb{U}}^{n}$ and $\bar{\hat{\mathbf{q}}}_{d}$ satisfies (11).
Proof.
Suppose that ${\hat{\mathbf{q}}}_{d}$ is a desired formation. Then it satisfies (3). Then the left of the $i$th equation of (11) satisfies
$d_{i}\hat{q}_{d_{i}}^{*}-\sum_{(i,j)\in E}\hat{q}_{d_{ij}}\hat{q}_{d_{j}}^{*}=d_{i}\hat{q}_{d_{i}}^{*}-\sum_{(i,j)\in E}\hat{q}_{d_{i}}^{*}\hat{q}_{d_{j}}\hat{q}_{d_{j}}^{*}=0.$
Thus, it satisfies (11). From this, we have (10).
On the other hand, suppose that $\hat{\mathbf{q}}\in{\hat{\mathbb{U}}^{n}}$ and satisfies (10). We now need to show that it satisfies (3). Then, by (10), we have ${\hat{L}}=\hat{Q}^{*}L\hat{Q}$ and
$\hat{q}_{d_{ij}}=\hat{q}_{i}^{*}\hat{q}_{j}.$
The sufficient condition holds from the definition.
Suppose that $\lambda$ is an eigenvalue of $L$ with an eigenvector $\mathbf{x}$. From the definition, we have
$L\mathbf{x}=\lambda\mathbf{x}\Leftrightarrow\hat{Q}\hat{L}\hat{Q}^{*}\mathbf{x}=\lambda\mathbf{x}\Leftrightarrow\hat{L}\hat{Q}^{*}\mathbf{x}=\hat{Q}^{*}\mathbf{x}\lambda\Leftrightarrow\hat{L}\hat{\mathbf{y}}=\hat{\mathbf{y}}\lambda,$
where $\hat{\mathbf{y}}=\hat{Q}^{*}\mathbf{x}$. In other words, all eigenvalues of $L$ are also right eigenvalues of $\hat{L}$ in the sense of [19]. By the spectral theory of $L$, the zero eigenvalue of $L$ has an eigenvector ${\bf 1}$. Therefore, zero is also a right eigenvalue of $\hat{L}$ and its corresponding eigenvector is $\hat{Q}^{*}{\bf 1}=\bar{\hat{\mathbf{q}}}_{d}$. By [19], a real right eigenvalue of a dual quaternion matrix can be simply called an eigenvalue of that matrix. Thus, zero is an eigenvalue of $\hat{L}$.
In addition, if $G$ has a spanning tree, then by Lemma 2.2, zero is a simple eigenvalue of $L$. In this case, the set of all vectors in $\hat{\mathbb{U}}^{n}$ that satisfy (11) is $\bar{\hat{\mathbf{q}}}_{d}\hat{c}$, where $\hat{c}\in\hat{\mathbb{U}}$. The sufficient condition holds from the fact that $\hat{c}^{*}\hat{\mathbf{q}}_{d}$ is also a desired formation.
This completes the proof. ∎
This theorem establishes the relation between desired formations and the null space of the dual quaternion Laplacian, and the relation between the dual quaternion Laplacian and the ordinary Laplacian of a UDQWDG $G$.
We should point out that (10) is only a necessary condition for $\Phi$ to be balanced. See Example 7.3 in [18] for a counterexample.
3 A Numerical Method to Determine if a Desired Relative Configuration Scheme is Reasonable or Not
Suppose that we have a connected directed graph $G=(V,E)$, and a desired relative configuration scheme $\left\{\hat{q}_{d_{ij}}:(i,j)\in E\right\}$. A practical question is: Whether such a desired relative configuration scheme is reasonable or not? Based upon the discussion in the last section, we have the following procedure to answer the above question.
1. If there are arcs $(i,j),(j,i)\in E$, check if (7) is satisfied or not. If (7) is not satisfied for such a pair of arcs, then the desired relative configuration scheme is not reasonable.
2. Construct the dual quaternion Laplacian matrix $\hat{L}$ by (32).
3. Solve
$\hat{L}{\hat{\mathbf{x}}}=\hat{\mathbf{0}},\text{ and }\hat{\mathbf{x}}\in\hat{\mathbb{U}}^{n}.$ (12)
If such a solution exists, then by Theorem 2.5, we check whether (10) holds. If so, it is a desired formation satisfying (3) and the desired relative configuration scheme is reasonable. Otherwise, it is not.
We may write $\hat{L}=L_{s}+L_{d}\epsilon$ and $\hat{\mathbf{x}}=\mathbf{x}_{s}+\mathbf{x}_{d}\epsilon$. Then (12) is equivalent to two quaternion equations
$L_{s}\mathbf{x}_{s}={\mathbf{0}},$ (13)
where $\mathbf{x}_{s}\in{\mathbb{U}}^{n}$, and
$L_{s}\mathbf{x}_{d}{=-L_{d}\mathbf{x}_{s}},$ (14)
where $x_{sj}x_{dj}^{*}+x_{dj}x_{sj}^{*}=0$ for $j=1,\dots,n$.
However, solving (13) directly may produce a zero solution which does not satisfy $\mathbf{x}_{s}\in{\mathbb{U}}^{n}$, even though the desired relative configuration scheme is reasonable. The situation may be even more complicated when the dimension of the null space of $L_{s}$ is greater than 1. In this case, we must exploit the null space of $L_{s}$ to see whether $\mathbf{x}_{s}\in{\mathbb{U}}^{n}$ exists. In the following, we make the following assumption.
Assumption 3.1.
Given a connected directed graph $G=(V,E)$, and a desired relative configuration scheme $\left\{\hat{q}_{d_{ij}}:(i,j)\in E\right\}$. Assume the standard part of the Laplacian matrix satisfies $\mathrm{rank}(L_{s})\geq n-1$.
Under Assumption 3.1, either (13) only has zero solution or all solutions of (13) may be denoted by
$[\mathbf{x}_{s}]=\{\mathbf{y}:\mathbf{y}=\mathbf{x}_{s}q,\ q\in\mathbb{Q}\},$ (15)
where $\mathbf{x}_{s}$ is a nonzero solution of (13). Similarly, when Assumption 3.1 holds and $\mathbf{x}_{s}$ satisfies (13), either (14) has no solution at all or all solutions of (14) may be denoted by
$[\mathbf{x}_{d}]=\{\mathbf{y}:\mathbf{y}=\mathbf{x}_{d}-\mathbf{x}_{s}q,\ q\in\mathbb{Q}\},$ (16)
where $\mathbf{x}_{d}$ is a solution of (14).
Different from the real cases, the nonzero solution of (13) is not unique even if the dimension of the null space of $L_{s}$ is equal to 1. In fact, for any $\mathbf{x}_{s}\in{\mathbb{U}}^{n}$ satisfying $L_{s}\mathbf{x}_{s}={\mathbf{0}}$ and any $q\in\mathbb{U}$, $\mathbf{x}_{s}q\in{\mathbb{U}}^{n}$ is also a solution of (13). Thus, we fix $x_{1s}=1$ and solve the reduced system
$L_{2s}\mathbf{x}_{2s}=-L_{1s},$ (17)
where $L_{1s}\in\mathbb{Q}^{n}$ and $L_{2s}\in\mathbb{Q}^{n\times(n-1)}$ are the first and remaining columns of $L_{s}$, respectively, and $\mathbf{x}_{2s}\in\mathbb{Q}^{n-1}$ is the last $n-1$ elements of $\mathbf{x}_{s}$. If (17) does not admit any solutions, then (13) only has zero solution and the configuration scheme is not reasonable. Otherwise, if (17) has a nonzero solution $\mathbf{x}_{s}$, we still have to check whether $\mathbf{x}_{s}\in{\mathbb{U}}^{n}$: If $\mathbf{x}_{s}\in{\mathbb{U}}^{n}$, then we go to next step and compute $\mathbf{x}_{d}$; Otherwise, by the expression in (15) and $x_{1s}=1$, we should choose $q\in\mathbb{U}$, which does not change the magnitude of the entries. Thus, we conclude that $\mathbf{y}\notin{\mathbb{U}}^{n}$ for any $\mathbf{y}\in[\mathbf{x}_{s}]$ and the configuration scheme is not reasonable.
Similarly, the solution of (14) is not unique when (13) has nonzero solutions. Thus, we fix $x_{1d}=0$ and solve the reduced system
$L_{2s}\mathbf{x}_{2d}=-L_{d}\mathbf{x}_{s},$ (18)
where $\mathbf{x}_{2d}\in\mathbb{Q}^{n-1}$ is the last $n-1$ entries of $\mathbf{x}_{d}$. In fact, for any solution $\mathbf{y}_{d}$ of (14), we can set $\mathbf{x}_{d}=\mathbf{y}_{d}-\mathbf{x}_{s}y_{1d}$ and derive $x_{1d}=0$. If (18) is solvable and $x_{sj}x_{dj}^{*}+x_{dj}x_{sj}^{*}=0$ for $j=1,\dots,n$, then the configuration scheme is reasonable. Otherwise, the configuration scheme is not reasonable since for any solution $\mathbf{y}\in[\mathbf{x}_{d}]$, there is
$x_{sj}y_{j}^{*}+y_{j}x_{sj}^{*}=x_{sj}x_{dj}^{*}+x_{dj}x_{sj}^{*}-x_{sj}(q+q^{*})x_{sj}^{*}=x_{sj}x_{dj}^{*}+x_{dj}x_{sj}^{*}-2q_{0},$
where $q_{0}$ is the standard part of $q$. This combining with $x_{1d}=0$ derives $q_{0}=0$. Thus we conclude that for any solution $\mathbf{y}\in[\mathbf{x}_{d}]$, there is $x_{sj}y_{j}^{*}+y_{j}x_{sj}^{*}=x_{sj}x_{dj}^{*}+x_{dj}x_{sj}^{*}$ for $i=1,\dots,n$ and the configuration scheme is not reasonable.
We summarize the whole procedure in the following algorithm.
Algorithm 1 A numerical method to determine if a desired relative configuration scheme is reasonable or not
1:a directed and connected graph $G=(V,E)$, and a desired relative configuration scheme $\left\{\hat{q}_{d_{ij}}:(i,j)\in E\right\}$.
2:If there are arcs $(i,j),(j,i)\in E$, check if (7) is satisfied or not. If (7) is not satisfied for such a pair of arcs, then Flag=0 and stop.
3:Construct the Laplacian matrix $L\in\mathbb{Q}^{n\times n}$ and solve (17).
4:if (17) is solvable and $\mathbf{x}_{s}\in\mathbb{U}^{n}$ then
5:      Solve (18).
6:      if (18) is solvable and $\hat{\mathbf{q}}_{d}\equiv\hat{\mathbf{x}}\in\hat{\mathbb{U}}^{n}$ satisfies (10then
7:           Flag=1 and stop.
8:      else
9:           Flag=0 and stop.
10:      end if
11:else
12:      Flag=0 and stop.
13:end if
14:Output: Flag.
Let us revisit the examples in Example 2.1. For the example of directed tree, we have
$L_{1s}^{(1)}=\begin{bmatrix}0\\ -q_{d_{21s}}\\ -q_{d_{31s}}\end{bmatrix},\quad L_{2s}^{(1)}=\begin{bmatrix}0&0\\ 1&0\\ 0&1\end{bmatrix},\text{ and }L_{d}^{(1)}=\begin{bmatrix}0&0&0\\ -q_{d_{21d}}&0&0\\ -q_{d_{31d}}&0&0\end{bmatrix}.$
Thus, by solving (17) and (18), we have $\mathbf{x}_{s}^{(1)}=[1,q_{d_{21s}},q_{d_{31s}}]^{\top}$ and $\mathbf{x}_{d}^{(1)}=[0,q_{d_{21d}},q_{d_{31d}}]^{\top}$, respectively. Namely, $\mathbf{x}^{(1)}=[1,\hat{q}_{d_{21}},\hat{q}_{d_{31}}]^{\top}\in\hat{\mathbb{U}}^{3}$ is a solution of (12) and the desired relative configuration scheme is reasonable for any $\hat{q}_{d_{21}},\hat{q}_{d_{31}}\in\hat{\mathbb{U}}$.
For the example of three points directed cycle, we have
$L_{1s}^{(2)}=\begin{bmatrix}1\\ -q_{d_{21s}}\\ 0\end{bmatrix},\quad L_{2s}^{(2)}=\begin{bmatrix}0&-q_{d_{13s}}\\ 1&0\\ -q_{d_{32s}}&1\end{bmatrix},\text{ and }L_{d}^{(2)}=\begin{bmatrix}0&0&-q_{d_{13d}}\\ -q_{d_{21d}}&0&0\\ 0&-q_{d_{32d}}&0\end{bmatrix}.$
Thus, system (17) is consistent only if $q_{d_{13s}}^{*}=q_{d_{32s}}q_{d_{21s}}$. In this case, we have $\mathbf{x}_{s}^{(2)}=[1,q_{d_{21s}},q_{d_{13s}}^{*}]^{\top}$. System (18) is consistent only if $q_{d_{13d}}^{*}=q_{d_{32s}}q_{d_{21d}}+q_{d_{32d}}q_{d_{21s}}$. In this case, we have $\mathbf{x}_{d}^{(2)}=[0,q_{d_{21d}},q_{d_{13d}}^{*}]^{\top}$. Namely, the desired relative configuration scheme is reasonable only if $\hat{q}_{d_{13}}^{*}=\hat{q}_{d_{32}}\hat{q}_{d_{21}}$, which has a solution $\mathbf{x}^{(2)}=[1,\hat{q}_{d_{21}},\hat{q}_{d_{13}}^{*}]^{\top}\in\hat{\mathbb{U}}^{3}$ for (12). We could verify that $\hat{\mathbf{q}}_{d}\equiv\hat{\mathbf{x}}^{(2)}$ satisfies (10).
In practice, we may convert (17) and (18) to real or complex equations, whose dimension is expanded to four times or twice of the original dimension. Then they are surely solvable. For instance, any quaternion linear equation $A\mathbf{x}=\mathbf{b}$ with $A\in\mathbb{Q}^{m\times n}$, $\mathbf{b}\in\mathbb{Q}^{m}$ can be equivalently rewritten to a real matrix equation
$\mathcal{R}(A)\mathcal{R}_{v}(\mathbf{x})=\mathcal{R}_{v}(\mathbf{b}),$ (19)
where $\mathcal{R}(\cdot)$ is a linear homeomorphic mapping from quaternion matrices (or vectors) to their real counterpart and $\mathcal{R}_{v}(\cdot)$ denotes the first column of $\mathcal{R}(\cdot)$. Specifically, let $A=[A_{0},A_{1},A_{2},A_{3}]$, then
$\mathcal{R}(A)=\begin{bmatrix}A_{0}&-A_{1}&-A_{2}&-A_{3}\\ A_{1}&A_{0}&-A_{3}&A_{2}\\ A_{2}&A_{3}&A_{0}&-A_{1}\\ A_{3}&-A_{2}&A_{1}&A_{0}\end{bmatrix}\text{ and }\mathcal{R}_{v}(A)=\begin{bmatrix}A_{0}\\ A_{1}\\ A_{2}\\ A_{3}\end{bmatrix}$
Several structure-preserving algorithms such as [5, 8, 11] can be used here to solve (19) efficiently.
4 Control Law
Now suppose that desired relative configuration scheme is reasonable. In this section, we propose the following control law
$\dot{\hat{\mathbf{z}}}=-\hat{K}\hat{L}\hat{\mathbf{z}},$ (20)
or equivalently,
$\dot{\hat{z}}_{i}=\hat{k}_{i}\left(\sum_{j\in\mathcal{N}_{i}}{\hat{q}_{d_{ij}}}\hat{z}_{j}-d_{i}\hat{z}_{i}\right)$ (21)
for $i=1,\dots,n$. Then the following theorem show the asymptotically convergence of our proposed control law. The Jordan canonical form is important in the proof.
Theorem 4.1.
Suppose that we have a directed graph $G=(V,E)$ that has a directed spanning tree and there is a reasonable desired relative configuration $\left\{\hat{q}_{d_{ij}}:(i,j)\in E\right\}$ on $G$. Let $\hat{\mathbf{q}}_{d}\in{\hat{\mathbb{U}}^{n}}$ be a desired formation satisfying (3). If there exists a positive diagonal matrix $K\in\mathbb{R}^{n\times n}$ such that $\hat{K}=\hat{Q}^{*}K\hat{Q}$, where $\hat{Q}=diag(\hat{\mathbf{q}}_{d})$, then the observed configuration vector $\hat{\mathbf{z}}=\hat{\mathbf{z}}(t)$ under control law (20) globally asymptotically converges to $\bar{\hat{\mathbf{q}}}_{d}\hat{c}$ with $\hat{c}\in{\hat{\mathbb{Q}}}$.
Proof.
From the assumptions we have
$\hat{K}\hat{L}=(\hat{Q}^{*}K\hat{Q})(\hat{Q}^{*}L\hat{Q})=\hat{Q}^{*}KL\hat{Q}.$
Since $K$ is a positive definite diagonal matrix, all eigenvalues of $KL$ have positive real parts except one zero eigenvalue. Denote the Jordan decomposition of $KL$ as $SJS^{-1}$. Without loss of generalization, let $J=diag(0,J_{2})$. Then $J_{2}$ is a Jordan matrix and all diagonal elements have positive real parts. Let $V=cS$. Then there is
$V^{-1}KLV=\begin{bmatrix}0&0\\ 0&J_{2}\end{bmatrix}.$ (22)
Here, $c=\frac{1}{{\bf 1}^{\top}s_{1}}$, $s_{1}$ is the first column of $S$. Therefore, it follows from the fact that the eigenvector corresponding to the zero eigenvalue of $L$ is $\mathbf{1}_{n}$ that the first column of $V\in\mathbb{R}^{n\times n}$ is $\mathbf{1}_{n}$. Define
${\hat{\mathbf{y}}}=\begin{bmatrix}\hat{y}_{1}\\ {\hat{\mathbf{y}}_{2}}\end{bmatrix}=V^{-1}\hat{Q}\hat{\mathbf{z}}.$
Then we have
$\begin{bmatrix}\dot{\hat{y}}_{1}\\ \dot{{\hat{\mathbf{y}}}}_{2}\end{bmatrix}=-\begin{bmatrix}0&0\\ 0&J_{2}\end{bmatrix}\begin{bmatrix}\hat{y}_{1}\\ {\hat{\mathbf{y}}_{2}}\end{bmatrix}.$ (23)
It follows that $\hat{y}_{1}(t)=\hat{y}_{1}(0)$ and ${\hat{\mathbf{y}}_{2}(t)=\exp(-J_{2}t)\hat{\mathbf{y}}(0)}$ asymptotically converges to a zero vector as $t\rightarrow\infty$. Therefore, $\hat{\mathbf{z}}(t)={\hat{Q}^{*}V\hat{\mathbf{y}}(t)}$ globally asymptotically converges to a formation $\bar{\hat{\mathbf{q}}}_{d}\hat{y}_{1}(0)$. This completes the proof. ∎
Now we draw a remark on the relationship between the control law in (20) and several state-of-the-art methods.
Remark 5.1. The graph Laplacian has a long history in the control law design, especially in the consensus problem. In 2004, Olfati-Saber and Murray [17] considered the linear system
$\dot{\mathbf{x}}(t)=-L\mathbf{x}(t),$ (24)
where $L$ is the graph Laplacian induced by the weighted digraph $G=(V,E,A)$ and $A$ is the weighted adjacency matrix. They showed that if $G$ is a connected undirected graph or a strongly connected digraph, then the algorithm in (24) asymptotically solves a consensus problem for all initial states. Furthermore, Olfati-Saber and Murray [17] established a direct connection between the algebraic connectivity of the network and the performance of the control. In 2005, Ren and Beard [22] considered consensus seeking in multi-agent systems under dynamically changing interaction topologies. In 2007, Olfati-Saber, Fax and Murray [16] provided a theoretical framework for analysis of consensus algorithms for multi-agent networked systems.
In 2020, Savino et al. [24] presented a solution based on dual quaternion algebra to the general problem of pose consensus for systems composed of multiple rigid-bodies, i.e.,
$\dot{\hat{\mathbf{y}}}=-L\hat{\mathbf{y}},$ (25)
where $\hat{\mathbf{y}}\in\hat{\mathbb{Q}}^{n}$ is a pure dual quaternion vector. They further showed the proposed pose consensus protocol has guaranteed convergence when the interaction among the agents is represented by directed graphs with directed spanning trees.
Our control law (20) can be regarded as a generalization of (24) in [17] and (25) in [24] from the consensus problem to a general formation control problem. In fact, for a consensus problem, we have $\hat{q}_{d_{1}}=\dots=\hat{q}_{d_{n}}$ and $\hat{q}_{d_{ij}}=1$ for all $(i,j)\in E$. Thus, there is $\hat{L}=L$. In this case, by setting $\hat{K}=I$, (20) reduces to $\dot{\hat{\mathbf{z}}}=-L\hat{\mathbf{z}}.$ This is exactly the control law (25) in [24].
5 Dual Quaternion Weighted Directed Graphs
The discussion in Sections 2-4 based on unit dual quaternions may be extended to a more general mathematical setting.
Suppose that we have a loopless directed graph $G=(V,E)$, a multiplication group $\Omega\subset\hat{\mathbb{Q}}$ and a weight function $\varphi:V\times V\to\hat{\mathbb{Q}}$. Here, $V$ is the vertex set and $E$ is the arc set. Note that $0\in\hat{\mathbb{Q}}$ and $1\in\Omega$.
Let $i,j\in V$. If $(i,j)\not\in E$, then we have ${\varphi}(i,j)=0$. If $(i,j)\in E$, then we have ${\varphi}(i,j)=\hat{q}_{ij}\in\Omega$. We call ${\varphi}(i,j)$ the weight of the arc $(i,j)$, and $\Omega$ the weight group of $\Phi$.
The triple $\Phi=(G,\Omega,\varphi)$ is called a dual quaternion weighted directed graph (DQWDG).
Example 1 Let $\Omega=\{1\}$. Then $\Phi$ is an ordinary graph.
Example 2 There is a rich literature on unit gain graphs. Examples of unit gain graphs include signed graphs [7], complex unit gain graphs [21], quaternion unit gain graphs [1] and dual quaternion unit gain graphs [4, 18]. Such unit gain graphs can be regarded as special cases of DQWDGs with two additional requirements: 1. For each $(i,j)\in E$, we have $(j,i)\in E$; 2. For each $(i,j)\in E$, we have $\hat{q}_{ji}={\hat{q}_{ij}^{*}}$. The adjacency and Laplacian matrices of a unit gain graph are Hermitian matrices, while the adjacency and Laplacian matrices of a DQWDG are not Hermitian matrices in general.
Example 3 If $\Omega=\{-1,1\}$, then we have signed directed graphs (SDG) [28] or directed signed graphs (DSG) [12]. Applications of SDG or DSG include process systems analysis [10, 29], complex network analysis [25] and data mining [9].
Example 4 If $\Omega={\mathbb{R}}_{++}$, the set of positive numbers, then we have nonnegative weighted directed graphs. Nonnegative weighted directed graphs have found broad applications in control engineering [6, 16].
Example 5 If $\Omega$ is the set of nonzero complex numbers, then we have complex WDGs [6]. Notably, complex WDGs are shown to be very useful in multi-agent formation control [13, 15]. Besides applications in formation control, networks with complex weights also arise in fields as diverse as quantum information, quantum chemistry, electrodynamics, rheology, and machine learning [2].
Example 6 If $\Omega=\hat{\mathbb{U}}$, then we have UDQWDG studied in the last three sections. The discussion in Sections 2 and 3 shows that UDQWDGs arise from applications. It also shows that there are some challenging issues there. For a UDQWDG, the adjacency matrix $A$ and the Laplacian matrix are not Hermitian matrices, while there is no theory on eigenvalues of non-Hermitian dual quaternion matrices.
These examples show that the class of DQWDGs is a general group of WDGs, which include all useful known WDGs. Thus, it is worth studying DQWDGs. In this section, we present a basic theory for DQWDGs.
A walk $W$ in a DQWDG $\Phi$ is defined in this way. It consists of vertices $i_{1},i_{2},\dots,i_{k}$. For $j=1,\dots,k-1$, either $(i_{j},i_{j+1})\in W$ or $(i_{j+1},i_{j})\in W$. If $(i_{j},i_{j+1})\in W$, we say that the arc $(i_{j},i_{j+1})$ is a forward arc of $W$ and let the shadow element $\hat{p}_{i_{j}i_{j+1}}=\hat{q}_{i_{j}i_{j+1}}$. If $(i_{j+1},i_{j})\in W$, we say that the arc $(i_{j+1},i_{j})$ is a backward arc of $W$ and let the shadow element $\hat{p}_{i_{j}i_{j+1}}=\hat{q}_{i_{j+1}i_{j}}^{-1}$. The weight of the walk $W$ is
$\varphi(W)=\prod_{j=1}^{k-1}\hat{p}_{i_{j}i_{j+1}}.$ (26)
The walk $W$ is neutral if $\varphi(W)$ is a real positive number. The DQWDG $\Phi$ is called balanced if every cycle $C\subseteq{E}$ is neutral.
The definition of balanced here is different from that of [6] since only forward arcs are considered in [6]. We exploit the difference by the following example, which is a slight modification of Example 4 of [6].
Example 5.1.
Consider two complex WDGs in Fig. 2. The complex WDG on the left is the same with Example 4 of [6]. It is unbalanced under our definition since $\phi(C)=-1$ for the cycle $C=\{1,2,3,1\}$. The complex WDG on the right is balanced.
We would like to point out that the complex WDG on the left of Fig. 2 is balanced under the definition of balanced in [6].
Figure 2: Two complex WDGs. Left: A unbalanced graph. Right: A balanced graph.
A potential function is a function $\theta:V\to\Omega$ such that for every arc $(i,j)\in E$,
$\varphi(i,j)=\theta(i)^{-1}\theta(j){c_{ij},}$ (27)
for a certain positive real number $c_{ij}$.
A switching function is a function $\zeta:V\to\Omega$ that switches the directed gain graph $\Phi=(G,\Omega,\varphi)$ to $\Phi^{\zeta}=(G,\Omega,\varphi^{\zeta})$, where
$\varphi^{\zeta}(i,j)=(\zeta(i))^{-1}\varphi(i,j)\zeta(j).$ (28)
Thus, in this case, $\Phi$ and $\Phi^{\zeta}$ are switching equivalent, denoted by $\Phi\sim\Phi^{\zeta}$. Denote $\bf 1$ as a weight function such that for any $(i,j)\in E$, we have ${\bf 1}(i,j)\equiv 1$. Then we have the following theorem.
Theorem 5.2.
Let $\Phi=(G,\Omega,\varphi)$ be a DQWDG, where $G=(V,E)$ is a connected directed graph. Then the following statements are equivalent.
(i) $\Phi$ is balanced.
(ii) $\Phi\sim(G,{\Omega,\varphi^{\zeta}})$, where ${\varphi^{\zeta}}(i,j)=|\varphi_{s}(i,j)|$ for all $(i,j)\in E$.
(iii) $\Phi$ has a potential function $\theta$.
Proof.
$(iii)\Rightarrow(ii)$ Suppose that there is a potential function $\theta:V\to\Omega$ and positive real numbers $c_{ij}$ such that $\varphi(i,j)=\theta(i)^{-1}\theta(j)c_{ij}$ for every arc $(i,j)\in E$. Since the dual part of $c_{ij}$ is zero, we have
$c_{ij}=\frac{|\varphi_{s}(i,j)||\theta_{s}(i)|}{|\theta_{s}(j)|},$ (29)
where $\varphi_{s}(i,j)$ and $\theta_{s}(i)$ are the standard parts of $\varphi(i,j)$ and $\theta(i)$, respectively. For $i=1,\dots,n$, let
$\zeta(i)={|\theta_{s}(i)|}\theta(i)^{-1}.$
Then we have
$\varphi^{\zeta}(i,j)={\frac{|\theta_{s}(j)|}{|\theta_{s}(i)|}}\theta(i)\varphi(i,j)\theta(j)^{-1}={\frac{|\theta_{s}(j)|}{|\theta_{s}(i)|}}c_{ij}=|\varphi_{s}(i,j)|.$ (30)
$(ii)\Rightarrow(i)$ It follows from (30) that $\varphi(i,j)=|\varphi_{s}(i,j)|\zeta(i)\zeta(j)^{-1}$. Given any cycle $C=\{i_{1},\dots,i_{k},i_{1}\}$, the weight of $C$ is equal to
$\varphi(C)=\hat{p}_{i_{1}i_{2}}\dots\hat{p}_{i_{k}i_{1}}=|\left(p_{i_{1}i_{2}}\right)_{s}|\dots|\left(p_{i_{k}i_{1}}\right)_{s}|,$
where $(p_{ij})_{s}$ is the standard part of the shadow element $\hat{p}_{ij}$. This implies that all cycles are neutral and $\Phi$ is balanced.
$(i)\Rightarrow(iii)$ Suppose that all cycles are neutral. Let $m$ be the number of edges of the underlying undirected graph of $G$. Since $G$ is connected, we have $m\geq n-1$. Let $M=m-n+1$. We now show by induction on $M$ that there is a potential function $\theta:V\to\Omega$ and positive real numbers $c_{ij}$ such that $\varphi(i,j)=\theta(i)^{-1}\theta(j)c_{ij}$ for every arc $(i,j)\in E$.
We first show this for $M=0,i.e.,m=n-1$. Then $G$ is a tree. Without loss of generality, we pick node $1$ as the root of this tree. Let $\theta(1)=1$. Then for all $j\not=1$, we may determine $c_{ij}$ by (29) and $\theta(j)$ by assigning $\theta(j)=\theta(i)\varphi(i,j)c_{ij}^{-1}$ from each father node $i$ to its children node $j$. When $M>0$, we may show the result by the induction approach similar to that of Theorem 2.4.
This completes the proof. ∎
Denote $\Phi=(G,\Omega,\varphi)$ as the complex WDG on the left of Fig. 2. Then We may check that $\theta(1)=1,\theta(2)=1,\theta(3)=\mathbf{i}$ is a potential function for $\Phi$ with $c_{21}=2$, $c_{12}=c_{13}=c_{23}=1$. Furthermore, let $\zeta(1)=\zeta(2)=1$ and $\zeta(3)=-\mathbf{i}$. Then $\varphi^{\zeta}(2,1)=2$ and $\varphi^{\zeta}(1,2)=\varphi^{\zeta}(1,3)=\varphi^{\zeta}(2,3)=1$.
We may also extend Theorem 2.5 to DQWDGs.
Let ${\hat{A}}=(\hat{a}_{ij})$, where $\hat{a}_{ij}={\varphi(i,j)}$ of $(i,j)\in V\times V$. The matrix ${\hat{A}}=({\hat{a}_{ij}})$ is called the adjacency matrix of $\Phi$.
The out-degree ${d_{i}}$ of a vertex $i$ is defined as
${d_{i}}=\sum_{j\in E_{i}}\left|{\varphi_{s}({i,j})}\right|,$ (31)
where $E_{i}=\left\{j:(i,j)\in E\right\}$ and $\varphi_{s}({i,j})$ is the standard part of $\varphi({i,j})$. Thus, the out-degree $d_{i}$ is a nonnegative real number. Note that by (1), this definition is consistent with the definition for UDQWDGs in Section 2.
The Laplacian matrix of a DQWDG $\Phi$ is defined by
$\hat{L}={D}-\hat{A},$ (32)
where $D$ is a diagonal matrix whose $i$-th diagonal element is equal to the out-degree $d_{i}$ of the $i$-th vertex.
Theorem 5.3.
Suppose that we have a DQWDG $\Phi=(G,\Omega,\varphi)$. Then $\Phi$ has a potential function $\theta$ if and only if there is
$\hat{\mathbf{\mathbf{y}}}=[\theta(1)^{-1}|\theta_{s}(1)|,\dots,\theta(n)^{-1}|\theta_{s}(n)|]^{\top}$ (33)
satisfies
$\hat{Y}^{-1}{\hat{L}}\hat{Y}=L=D-A,$ (34)
where $\hat{Y}=\mathrm{diag}(\hat{\mathbf{\mathbf{y}}})$, $A=(a_{ij})$ is the weighted adjacency matrix with $a_{ij}=|\varphi_{s}(i,j)|$ for all $(i,j)\in{V\times V}$.
Furthermore, assume that $\Phi$ has a potential function $\theta$. Then all eigenvalues of $L$ are right eigenvalues of $\hat{L}$, which are complex numbers, and zero is an eigenvalue of $\hat{L}$ that satisfies
$\hat{L}\hat{\mathbf{\mathbf{y}}}=\hat{\mathbf{0}}.$ (35)
In other words, all inverse potential vectors are eigenvectors of $\hat{L}$, corresponding to the zero eigenvalue of $\hat{L}$.
If, in addition, $G$ has a directed spanning tree, then $\theta$ is a potential function if and only if it is invertible and $\hat{\mathbf{\mathbf{y}}}$ defined by (33) satisfies $|\mathbf{y}_{s}|={\bf 1}$ and (35).
Proof.
Suppose that $\Phi$ has a potential function $\theta$ which satisfies (27). Then for all $(i,j)\in E$, we have
$\hat{y}_{i}^{-1}\varphi(i,j)\hat{y}_{j}={\frac{|\theta_{s}(j)|}{|\theta_{s}(i)|}}\theta(i)\theta(i)^{-1}\theta(j)c_{ij}\theta(j)^{-1}=|\varphi_{s}(i,j)|,$
where $\hat{\mathbf{\mathbf{y}}}$ is defined by (33), and the last equality follows from (29). For the diagonal elements, we have
$\hat{y}_{i}^{-1}d_{i}\hat{y}_{i}=d_{i},$
where the equality follows from the fact that $d_{i}$ is a real number, which is commutative with any dual quaternion numbers. This derives (34).
On the other hand, suppose (34) holds. Then we could verify that (27) holds with $c_{ij}$ satisfies (29).
Suppose that $\lambda$ is an eigenvalue of $L$ with an eigenvector $\mathbf{x}$. From the definition, we have
$L\mathbf{x}=\lambda\mathbf{x}\Longleftrightarrow\hat{Y}^{-1}{\hat{L}}\hat{Y}\mathbf{x}=\lambda\mathbf{x}\Longleftrightarrow\hat{L}\hat{Y}\mathbf{x}=\hat{Y}\mathbf{x}\lambda\Longleftrightarrow\hat{L}\hat{\mathbf{z}}=\hat{\mathbf{z}}\lambda,$
where $\hat{\mathbf{z}}=\hat{Y}\mathbf{x}$. In other words, all eigenvalues of $L$ are also right eigenvalues of $\hat{L}$ in the sense of [19]. By the spectral theory of $L$, the zero eigenvalue of $L$ has an eigenvector ${\bf 1}$. Therefore, zero is also a right eigenvalue of $\hat{L}$ and its corresponding eigenvector is $\hat{\mathbf{y}}$. By [19], a real right eigenvalue of a dual quaternion matrix can be simply called an eigenvalue of that matrix. Thus, zero is an eigenvalue of $\hat{L}$.
In addition, if $G$ has a directed spanning tree, then by Lemma 2.2, zero is a simple eigenvalue of $L$. In this case, the set of all invertible vectors that satisfy (35) and $|\mathbf{y}_{s}|={\bf 1}$ is $[\hat{\mathbf{y}}]=\hat{\mathbf{y}}\hat{c}$, where $\hat{c}\in\hat{\mathbb{U}}$. By (33), we could construct the set $[\theta]=\hat{c}^{-1}\theta$ corresponding to $[\hat{\mathbf{y}}]$. The sufficient condition holds from the fact that $\hat{c}^{-1}\theta$ is also a potential function.
This completes the proof. ∎
In [6], Dong and Lin studied complex WDGs and derived necessary and sufficient conditions for the Laplacian matrix to be singular/ nonsingular. The contribution of Dong and Lin is that their results were established without restriction that the adjacency matrix is Hermitian. As we said in this section, DQWDG is a more general mathematical setting, and there is no theory on eigenvalues of non-Hermitian dual quaternion matrices. Hence, it would be interesting to extend the results of Dong and Lin to DQWDGs. We leave this as a future research work.
Acknowledgment This work was partially supported by the R&D project of Pazhou Lab (Huangpu) (Grant no. 2023K0603), the National Natural Science Foundation of China (No. 12131004), and the Fundamental Research Funds for the Central Universities (Grant No. YWF-22-T-204).
Data availability Data will be made available on reasonable request.
Conflict of interest The authors declare no Conflict of interest.